Screen

The virtual terminal manager

for Version 4.6.2 Feb 2017 Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-2017 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

Short Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Invoking Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Customizing Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 5 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6 New Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7 Selecting a Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8 Session Management Commands . . . . . . . . . . . . . . . . . . . . . . . . 33 9 Regions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 10 Window Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 11 Virtual Terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 12 Copy and Paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 13 Subprocess Execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 14 Key Binding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 15 Flow Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 16 Termcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 17 The Message Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 18 Logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 19 Startup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 20 Miscellaneous commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 21 String Escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 22 Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 23 Files Referenced. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 24 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 25 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 26 Installation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 Keystroke Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Getting Started. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Invoking Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Customizing Screen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.1 The .screenrc file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2 Source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3 Colon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5.1 Default Key Bindings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.2 Command Summary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

6 New Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

6.1 Chdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.2 Screen Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.3 Setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.4 Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.5 Term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.6 Window Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.7 Window Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

7 Selecting a Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

7.1 Moving Back and Forth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.2 Other Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.3 Select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.4 Windowlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

8 Session Management Commands . . . . . . . . . . . . . 33

8.1 Detach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 8.2 Power Detach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 8.3 Lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.4 Multiuser Session . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

8.4.1 Multiuser. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.4.2 Acladd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.4.3 Aclchg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 8.4.4 Acldel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 8.4.5 Aclgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 8.4.6 Displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 8.4.7 aclumask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 8.4.8 Wall. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 8.4.9 Writelock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 8.4.10 Su . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

8.5 Session Name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 8.6 Suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 8.7 Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

9 Regions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

9.1 Split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 9.2 Focus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 9.3 Only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 9.4 Remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 9.5 Resize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 9.6 Caption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 9.7 Fit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 9.8 Focusminsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 9.9 Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

10 Window Settings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

10.1 Naming Windows (Titles) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 10.1.1 Title Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 10.1.2 Dynamic Titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 10.1.3 Setting up your prompt for shell titles . . . . . . . . . . . . . . . . . . . 46 10.1.4 Setting up shell titles in your .screenrc . . . . . . . . . . . . . . . . . 46

10.2 Console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 10.3 Kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 10.4 Login . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 10.5 Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 10.6 Monitoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 10.7 Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 10.8 Hardstatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 10.9 Mousetrack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

11 Virtual Terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

11.1 Control Sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 11.2 Input Translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 11.3 Digraph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 11.4 Bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 11.5 Clear. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 11.6 Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 11.7 Redisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 11.8 Wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 11.9 Reset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 11.10 Window Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 11.11 Character Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

12 Copy and Paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

12.1 Copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.1.1 CR/LF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.1.2 Scrollback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.1.3 Markkeys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.1.4 Movement Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 12.1.5 Marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 12.1.6 Repeat Count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 12.1.7 Searching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 12.1.8 Specials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

12.2 Paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 12.3 Registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 12.4 Screen Exchange . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 12.5 History. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

13 Subprocess Execution . . . . . . . . . . . . . . . . . . . . . . . . 71

13.1 Exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 13.2 Using Exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

14 Key Binding. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

14.1 The bind command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 14.2 Examples of the bind command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 14.3 Command Character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 14.4 Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 14.5 Bindkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 14.6 Bindkey Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 14.7 Bindkey Control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

15 Flow Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

15.1 About screen flow control settings . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 15.2 Flow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 15.3 XON and XOFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

16 Termcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

16.1 Choosing the termcap entry for a window . . . . . . . . . . . . . . . . . . . . . 79 16.2 Write out the window’s termcap entry . . . . . . . . . . . . . . . . . . . . . . . . 80 16.3 The termcap command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 16.4 Termcap Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 16.5 Special Terminal Capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 16.6 Autonuke . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 16.7 Obuflimit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 16.8 Character Translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

17 The Message Line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

17.1 Using the message line from your program . . . . . . . . . . . . . . . . . . . . 85 17.2 Hardware Status Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.3 Display Last Message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 17.4 Message Wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

18 Logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

18.1 hardcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 18.2 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

19 Startup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

19.1 echo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 19.2 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 19.3 Startup Message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

20 Miscellaneous commands . . . . . . . . . . . . . . . . . . . . . 91

20.1 At . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 20.2 Break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 20.3 Bumpleft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 20.4 Bumpright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 20.5 Collapse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 20.6 Debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 20.7 License. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 20.8 Nethack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 20.9 Nonblock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 20.10 Number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 20.11 Time. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 20.12 Verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 20.13 Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 20.14 Zombie. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 20.15 Printcmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 20.16 Rendition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 20.17 Sorendition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 20.18 Attrcolor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 20.19 Setsid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 20.20 Eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 20.21 Maxwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 20.22 Backtick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 20.23 Screen Saver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 20.24 Zmodem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

21 String Escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

22 Environment Variables . . . . . . . . . . . . . . . . . . . . . . 103 23 Files Referenced . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

24 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

25 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

25.1 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 25.2 Reporting Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 25.3 Availability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

26 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

26.1 Socket Directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 26.2 Compiling Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

Keystroke Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

1 Overview

Screen is a full-screen window manager that multiplexes a physical terminal between several processes, typically interactive shells. Each virtual terminal provides the functions of the DEC VT100 terminal and, in addition, several control functions from the ISO 6429 (ECMA 48, ANSI X3.64) and ISO 2022 standards (e.g. insert/delete line and support for multiple character sets). There is a scrollback history buffer for each virtual terminal and a copyand-paste mechanism that allows the user to move text regions between windows.

When screen is called, it creates a single window with a shell in it (or the specified command) and then gets out of your way so that you can use the program as you normally would. Then, at any time, you can create new (full-screen) windows with other programs in them (including more shells), kill the current window, view a list of the active windows, turn output logging on and off, copy text between windows, view the scrollback history, switch between windows, etc. All windows run their programs completely independent of each other. Programs continue to run when their window is currently not visible and even when the whole screen session is detached from the user’s terminal.

When a program terminates, screen (per default) kills the window that contained it. If this window was in the foreground, the display switches to the previously displayed window; if none are left, screen exits. Shells usually distinguish between running as loginshell or sub-shell. Screen runs them as sub-shells, unless told otherwise (See shell .screenrc command).

Everything you type is sent to the program running in the current window. The only exception to this is the one keystroke that is used to initiate a command to the window manager. By default, each command begins with a control-a (abbreviated C-a from now on), and is followed by one other keystroke. The command character (see Section 14.3 [Command Character], page 74) and all the key bindings (see Chapter 14 [Key Binding], page 73) can be fully customized to be anything you like, though they are always two characters in length.

Screen does not understand the prefix C- to mean control, although this notation is used in this manual for readability. Please use the caret notation (^A instead of C-a) as arguments to e.g. the escape command or the -e option. Screen will also print out control characters in caret notation.

The standard way to create a new window is to type C-a c. This creates a new window running a shell and switches to that window immediately, regardless of the state of the process running in the current window. Similarly, you can create a new window with a custom command in it by first binding the command to a keystroke (in your .screenrc file or at the C-a : command line) and then using it just like the C-a c command. In addition, new windows can be created by running a command like:

     screen emacs prog.c

from a shell prompt within a previously created window. This will not run another copy of screen, but will instead supply the command name and its arguments to the window manager (specified in the $STY environment variable) who will use it to create the new window. The above example would start the emacs editor (editing prog.c) and switch to its window. - Note that you cannot transport environment variables from the invoking shell to the application (emacs in this case), because it is forked from the parent screen process, not from the invoking shell.

If /etc/utmp is writable by screen, an appropriate record will be written to this file for each window, and removed when the window is closed. This is useful for working with talk, script, shutdown, rsend, sccs and other similar programs that use the utmp file to determine who you are. As long as screen is active on your terminal, the terminal’s own record is removed from the utmp file. See Section 10.4 [Login], page 47.

2 Getting Started

Before you begin to use screen you’ll need to make sure you have correctly selected your terminal type, just as you would for any other termcap/terminfo program. (You can do this by using tset, qterm, or just set term=mytermtype, for example.)

If you’re impatient and want to get started without doing a lot more reading, you should remember this one command: C-a ? (see Chapter 14 [Key Binding], page 73). Typing these two characters will display a list of the available screen commands and their bindings. Each keystroke is discussed in the section on keystrokes (see Section 5.1 [Default Key Bindings], page 11). Another section (see Chapter 4 [Customization], page 9) deals with the contents

of your .screenrc.

If your terminal is a “true” auto-margin terminal (it doesn’t allow the last position on the screen to be updated without scrolling the screen) consider using a version of your terminal’s termcap that has automatic margins turned off. This will ensure an accurate and optimal update of the screen in all circumstances. Most terminals nowadays have “magic” margins (automatic margins plus usable last column). This is the VT100 style type and perfectly suited for screen. If all you’ve got is a “true” auto-margin terminal screen will be content to use it, but updating a character put into the last position on the screen may not be possible until the screen scrolls or the character is moved into a safe position in some other way. This delay can be shortened by using a terminal with insert-character capability.

See Section 16.5 [Special Capabilities], page 82, for more information about telling screen what kind of terminal you have.

3 Invoking Screen

Screen has the following command-line options:

‘-a’ Include all capabilities (with some minor exceptions) in each window’s termcap, even if screen must redraw parts of the display in order to implement a function.

‘-A’ Adapt the sizes of all windows to the size of the display. By default, screen may try to restore its old window sizes when attaching to resizable terminals (those with ‘WS’ in their descriptions, e.g. suncmd or some varieties of xterm).

‘-c file’ Use file as the user’s configuration file instead of the default of

$HOME/.screenrc.
‘-d [pid.sessionname]’
‘-D [pid.sessionname]’

Do not start screen, but instead detach a screen session running elsewhere (see Section 8.1 [Detach], page 33). ‘-d’ has the same effect as typing C-a d from the controlling terminal for the session. ‘-D’ is the equivalent to the power detach key. If no session can be detached, this option is ignored. In combination with the -r/-R option more powerful effects can be achieved:

-d -r Reattach a session and if necessary detach it first.

-d -R Reattach a session and if necessary detach or even create it first.

-d -RR Reattach a session and if necessary detach or create it. Use the first session if more than one session is available.

-D -r Reattach a session. If necessary detach and logout remotely first.

-D -R Attach here and now. In detail this means: If a session is running, then reattach. If necessary detach and logout remotely first. If it was not running create it and notify the user. This is the author’s favorite.

-D -RR Attach here and now. Whatever that means, just do it.

Note: It is a good idea to check the status of your sessions with screen -list before using this option.

‘-e xy’ Set the command character to x, and the character generating a literal command character (when typed after the command character) to y. The defaults are C-a and a, which can be specified as ‘-e^Aa’. When creating a screen session, this option sets the default command character. In a multiuser session all users added will start off with this command character. But when attaching to an already running session, this option only changes the command character of the attaching user. This option is equivalent to the commands defescape or escape respectively. (see Section 14.3 [Command Character], page 74).

‘-f’
‘-fn’

‘-fa’ Set flow-control to on, off, or automatic switching mode, respectively. This option is equivalent to the defflow command (see Chapter 15 [Flow Control], page 77).

‘-h num’ Set the history scrollback buffer to be num lines high. Equivalent to the defscrollback command (see Section 12.1 [Copy], page 65).

‘-i’ Cause the interrupt key (usually C-c) to interrupt the display immediately when flow control is on. This option is equivalent to the interrupt argument to the defflow command (see Chapter 15 [Flow Control], page 77). Its use is discouraged.

‘-l’

‘-ln’ Turn login mode on or off (for /etc/utmp updating). This option is equivalent to the deflogin command (see Section 10.4 [Login], page 47).

‘-ls [match]’
‘-list [match]’

Do not start screen, but instead print a list of session identification strings (usually of the form pid.tty.host; see Section 8.5 [Session Name], page 37). Sessions marked ‘detached’ can be resumed with screen -r. Those marked ‘attached’ are running and have a controlling terminal. If the session runs in multiuser mode, it is marked ‘multi’. Sessions marked as ‘unreachable’ either live on a different host or are dead. An unreachable session is considered dead, when its name matches either the name of the local host, or the specified parameter, if any. See the -r flag for a description how to construct matches. Sessions marked as ‘dead’ should be thoroughly checked and removed. Ask your system administrator if you are not sure. Remove sessions with the ‘-wipe’ option.

‘-L’ Tell screen to turn on automatic output logging for the windows.

‘-Logfile "file"’

By default logfile name is "screenlog.0". You can set new logfile name with the

-Logfile option.

‘-m’ Tell screen to ignore the $STY environment variable. When this option is used, a new session will always be created, regardless of whether screen is being called from within another screen session or not. This flag has a special meaning in connection with the ‘-d’ option:

-d -m Start screen in detached mode. This creates a new session but doesn’t attach to it. This is useful for system startup scripts.

-D -m This also starts screen in detached mode, but doesn’t fork a new process. The command exits if the session terminates.

‘-O’ Select a more optimal output mode for your terminal rather than true VT100 emulation (only affects auto-margin terminals without ‘LP’). This can also be set in your .screenrc by specifying ‘OP’ in the termcap command.

‘-p name_or_number|-|=|+’

Preselect a window. This is useful when you want to reattach to a specific window or you want to send a command via the ‘-X’ option to a specific window. As with screen’s select command, ‘-’ selects the blank window. As a special case for reattach, ‘=’ brings up the windowlist on the blank window, while a ‘+’ will create new window. The command will not be executed if the specified window could not be found.

‘-q’ Suppress printing of error messages. In combination with ‘-ls’ the exit value is set as follows: 9 indicates a directory without sessions. 10 indicates a directory with running but not attachable sessions. 11 (or more) indicates 1 (or more) usable sessions. In combination with ‘-r’ the exit value is as follows: 10 indicates that there is no session to resume. 12 (or more) indicates that there are 2 (or more) sessions to resume and you should specify which one to choose. In all other cases ‘-q’ has no effect.

‘-Q’ Some commands now can be queried from a remote session using this flag, e.g. ’screen -Q windows’. The commands will send the response to the stdout of the querying process. If there was an error in the command, then the querying process will exit with a non-zero status. The commands that can be queried now are: echoinfolastmsgnumberselect

time title windows
‘-r [pid.sessionname]’
‘-r sessionowner/[pid.sessionname]’

Resume a detached screen session. No other options (except combinations with ‘-d’ or ‘-D’) may be specified, though the session name (see Section 8.5 [Session Name], page 37) may be needed to distinguish between multiple detached screen sessions. The second form is used to connect to another user’s screen session which runs in multiuser mode. This indicates that screen should look for sessions in another user’s directory. This requires setuid-root.

‘-R’ resumes screen only when it’s unambiguous which one to attach, usually when only one screen is detached. Otherwise lists available sessions.

‘-RR’ Resume the first appropriate detached screen session. If successful, all other command-line options are ignored. If no detached session exists, start a new session using the specified options, just as if ‘-R’ had not been specified. This option is set by default if screen is run as a login-shell (actually screen uses ‘-xRR’ in that case). For combinations with the ‘-D’/‘-d’ option see there.

‘-s program’

Set the default shell to be program. By default, screen uses the value of the environment variable $SHELL, or /bin/sh if it is not defined. This option is equivalent to the shell command (see Section 6.4 [Shell], page 28). See also there.

‘-S sessionname’

Set the name of the new session to sessionname. This option can be used to specify a meaningful name for the session in place of the default tty.host suffix. This name identifies the session for the screen -list and screen -r commands. This option is equivalent to the sessionname command (see Section 8.5 [Session Name], page 37).

‘-t name’ Set the title (name) for the default shell or specified program. This option is equivalent to the shelltitle command (see Section 6.4 [Shell], page 28).

‘-T term’ Set the $TERM enviroment variable using the specified term as opposed to the default setting of screen.

‘-U’ Run screen in UTF-8 mode. This option tells screen that your terminal sends and understands UTF-8 encoded characters. It also sets the default encoding for new windows to ‘utf8’.

‘-v’ Print the version number.

‘-wipe [match]’

List available screens like screen -ls, but remove destroyed sessions instead of marking them as ‘dead’. An unreachable session is considered dead, when its name matches either the name of the local host, or the explicitly given parameter, if any. See the -r flag for a description how to construct matches.

‘-x’ Attach to a session which is already attached elsewhere (multi-display mode). Screen refuses to attach from within itself. But when cascading multiple screens, loops are not detected; take care.

‘-X’ Send the specified command to a running screen session. You may use the -S option to specify the screen session if you have several running. You can use the -d or -r option to tell screen to look only for attached or detached screen sessions. Note that this command doesn’t work if the session is password protected.

4 Customizing Screen

You can modify the default settings for screen to fit your tastes either through a personal .screenrc file which contains commands to be executed at startup, or on the fly using the colon command.

4.1 The .screenrc file

When screen is invoked, it executes initialization commands from the files .screenrc in the user’s home directory and /usr/local/etc/screenrc. These defaults can be overridden in the following ways: For the global screenrc file screen searches for the environment variable $SYSSCREENRC (this override feature may be disabled at compile-time). The user specific screenrc file is searched for in $SCREENRC, then $HOME/.screenrc. The command line option ‘-c’ specifies which file to use (see Chapter 3 [Invoking Screen], page 5. Commands in these files are used to set options, bind commands to keys, and to automatically establish one or more windows at the beginning of your screen session. Commands are listed one per line, with empty lines being ignored. A command’s arguments are separated by tabs or spaces, and may be surrounded by single or double quotes. A ‘#’ turns the rest of the line into a comment, except in quotes. Unintelligible lines are warned about and ignored. Commands may contain references to environment variables. The syntax is the shell-like $VAR or ${VAR}. Note that this causes incompatibility with previous screen versions, as now the ’$’-character has to be protected with ’\’ if no variable substitution is intended. A string in single-quotes is also protected from variable substitution.

Two configuration files are shipped as examples with your screen distribution: etc/screenrc and etc/etcscreenrc. They contain a number of useful examples for various commands.

4.2 Source

[Command] source file

(none)

Read and execute commands from file file. Source commands may be nested to a maximum recursion level of ten. If file is not an absolute path and screen is already processing a source command, the parent directory of the running source command file is used to search for the new command file before screen’s current directory. Note that termcap/terminfo/termcapinfo commands only work at startup and reattach time, so they must be reached via the default screenrc files to have an effect.

4.3 Colon

Customization can also be done online, with this command:

[Command] colon

     (C-a :)

Allows you to enter .screenrc command lines. Useful for on-the-fly modification of key bindings, specific window creation and changing settings. Note that the set keyword no longer exists, as of version 3.3. Change default settings with commands 10 Screen User’s Manual

starting with ‘def’. You might think of this as the ex command mode of screen, with copy as its vi command mode (see Chapter 12 [Copy and Paste], page 65).

5 Commands

A command in screen can either be bound to a key, invoked from a screenrc file, or called from the colon prompt (see Chapter 4 [Customization], page 9). As of version 3.3, all commands can be bound to keys, although some may be less useful than others. For a number of real life working examples of the most important commands see the files etc/screenrc and etc/etcscreenrc of your screen distribution.

In this manual, a command definition looks like this:

– Command: command [-n] ARG1 [ARG2] … (keybindings) This command does something, but I can’t remember what.

An argument in square brackets (‘[]’) is optional. Many commands take an argument of ‘on’ or ‘off’, which is indicated as state in the definition.

5.1 Default Key Bindings

As mentioned previously, each keyboard command consists of a C-a followed by one other character. For your convenience, all commands that are bound to lower-case letters are also bound to their control character counterparts (with the exception of C-a a; see below). Thus, both C-a c and C-a C-c can be used to create a window.

The following table shows the default key bindings:

C-a ' (select)

Prompt for a window identifier and switch. See Chapter 7 [Selecting], page 31.

C-a " (windowlist -b)

Present a list of all windows for selection. See Chapter 7 [Selecting], page 31.

C-a 0...9, -

(select 0…select 9, select -)

Switch to window number 0…9, or the blank window. See Chapter 7 [Selecting], page 31.

C-a Tab (focus)

Switch the input focus to the next region. See Chapter 9 [Regions], page 39.

C-a C-a (other)

Toggle to the window displayed previously. If this window does no longer exist, other has the same effect as next. See Chapter 7 [Selecting], page 31.

C-a a (meta)

Send the command character (C-a) to window. See escape command. See Section 14.3 [Command Character], page 74.

C-a A (title)

Allow the user to enter a title for the current window. See Section 10.1 [Naming Windows], page 45.

C-a b

C-a C-b (break)

Send a break to the tty. See Section 20.2 [Break], page 91.

C-a B (pow break)

Close and reopen the tty-line. See Section 20.2 [Break], page 91.

C-a c

C-a C-c (screen)

Create a new window with a shell and switch to that window. See Section 6.2 [Screen Command], page 27.

C-a C (clear)

Clear the screen. See Section 11.5 [Clear], page 60.

C-a d

C-a C-d (detach)

Detach screen from this terminal. See Section 8.1 [Detach], page 33.

C-a D D (pow detach)

Detach and logout. See Section 8.2 [Power Detach], page 33.

C-a f
C-a C-f (flow)

Cycle flow among ‘on’, ‘off’ or ‘auto’. See Section 15.2 [Flow], page 77.

C-a F  (fit)

Resize the window to the current region size. See Section 9.7 [Fit], page 41.

C-a C-g (vbell)

Toggle visual bell mode. See Section 11.4 [Bell], page 59.

C-a h (hardcopy)

Write a hardcopy of the current window to the file “hardcopy.n”. See Section 18.1 [Hardcopy], page 87.

C-a H (log)

Toggle logging of the current window to the file “screenlog.n”. See Section 18.2 [Log], page 87.

C-a i
C-a C-i (info)

Show info about the current window. See Section 11.6 [Info], page 60.

C-a k
C-a C-k (kill)

Destroy the current window. See Section 10.3 [Kill], page 47.

C-a l

C-a C-l (redisplay)

Fully refresh the current window. See Section 11.7 [Redisplay], page 61.

C-a L (login)

Toggle the current window’s login state. See Section 10.4 [Login], page 47.

C-a m

C-a C-m (lastmsg)

Repeat the last message displayed in the message line. See Section 17.3 [Last Message], page 85.

C-a M (monitor) Toggle monitoring of the current window. See Section 10.6 [Monitor], page 48.

C-a SPC
C-a n
C-a C-n (next)

Switch to the next window. See Chapter 7 [Selecting], page 31.

C-a N (number)

Show the number (and title) of the current window. See Section 20.10 [Number], page 93.

C-a p
C-a C-p
C-a C-h
C-a BackSpace

(prev)

Switch to the previous window (opposite of C-a n). See Chapter 7 [Selecting], page 31.

C-a q
C-a C-q (xon)

Send a ^Q (ASCII XON) to the current window. See Section 15.3 [XON/XOFF], page 78.

C-a Q (only)

Delete all regions but the current one. See Chapter 9 [Regions], page 39.

C-a r
C-a C-r (wrap)

Toggle the current window’s line-wrap setting (turn the current window’s automatic margins on or off). See Section 11.8 [Wrap], page 61.

C-a s
C-a C-s (xoff)

Send a ^S (ASCII XOFF) to the current window. See Section 15.3 [XON/XOFF], page 78.

C-a S (split)

Split the current region horizontally into two new ones. See Chapter 9 [Regions], page 39.

C-a t
C-a C-t (time)

Show the load average and xref. See Section 20.11 [Time], page 93.

C-a v (version)

Display the version and compilation date. See Section 20.13 [Version], page 94.

C-a C-v (digraph)

Enter digraph. See Section 11.3 [Digraph], page 55.

C-a w

C-a C-w (windows)

Show a list of active windows. See Section 10.7 [Windows], page 49.

C-a W (width)

Toggle between 80 and 132 columns. See Section 11.10 [Window Size], page 62.

C-a x

C-a C-x (lockscreen)

Lock your terminal. See Section 8.3 [Lock], page 34.

C-a X (remove)

Kill the current region. See Chapter 9 [Regions], page 39.

C-a z

C-a C-z (suspend)

Suspend screen. See Section 8.6 [Suspend], page 38.

C-a Z (reset)

Reset the virtual terminal to its “power-on” values. See Section 11.9 [Reset], page 62.

C-a . (dumptermcap)

Write out a .termcap file. See Section 16.2 [Dump Termcap], page 80.

C-a ? (help)

Show key bindings. See Section 14.4 [Help], page 74.

C-a \ (quit)

Kill all windows and terminate screen. See Section 8.7 [Quit], page 38.

C-a : (colon)

Enter a command line. See Section 4.3 [Colon], page 9.

C-a [
C-a C-[
C-a ESC (copy)

Enter copy/scrollback mode. See Section 12.1 [Copy], page 65.

C-a ]

C-a C-] (paste .)

Write the contents of the paste buffer to the stdin queue of the current window. See Section 12.2 [Paste], page 68.

C-a {

C-a } (history)

Copy and paste a previous (command) line. See Section 12.5 [History], page 70.

C-a > (writebuf)

Write the paste buffer out to the screen-exchange file. See Section 12.4 [Screen Exchange], page 69.

C-a < (readbuf)

Read the screen-exchange file into the paste buffer. See Section 12.4 [Screen Exchange], page 69.

C-a = (removebuf)

Delete the screen-exchange file. See Section 12.4 [Screen Exchange], page 69.

C-a _(silence)

Start/stop monitoring the current window for inactivity. See Section 10.6 [Monitor], page 48.

C-a | (split -v)

Split the current region vertically into two new ones. See Chapter 9 [Regions], page 39.

C-a , (license)

Show the copyright page. See Section 20.7 [License], page 92.

C-a * (displays)

Show the listing of attached displays. See Section 8.4.6 [Displays], page 35.

5.2 Command Summary

acladd usernames

Allow other users in this session. See Section 8.4 [Multiuser Session], page 34.

aclchg usernames permbits list

Change a user’s permissions. See Section 8.4 [Multiuser Session], page 34.

acldel username

Disallow other user in this session. See Section 8.4 [Multiuser Session], page 34.

aclgrp usrname [groupname]

Inherit permissions granted to a group leader. See Section 8.4 [Multiuser Session], page 34.

aclumask [users]+/-bits ...

Predefine access to new windows. See Section 8.4.7 [Umask], page 36.

activity message

Set the activity notification message. See Section 10.6 [Monitor], page 48.

addacl usernames

Synonym to acladd. See Section 8.4 [Multiuser Session], page 34.

allpartial state

Set all windows to partial refresh. See Section 11.7 [Redisplay], page 61.

altscreen state

Enables support for the "alternate screen" terminal capability. See Section 11.7 [Redisplay], page 61.

at [ident][#|*|%] command [args]

Execute a command at other displays or windows. See Section 20.1 [At], page 91.

attrcolor attrib [attribute/color-modifier]

Map attributes to colors. See Section 20.18 [Attrcolor], page 95.

autodetach state

Automatically detach the session on SIGHUP. See Section 8.1 [Detach], page 33.

autonuke state

Enable a clear screen to discard unwritten output. See Section 16.6 [Autonuke], page 83.

backtick id lifespan autorefresh command [args]

Define a command for the backtick string escape. See Section 20.22 [Backtick], page 96.

bce [state]

Change background color erase. See Section 11.11 [Character Processing], page 62.

bell_msg [message]

Set the bell notification message. See Section 11.4 [Bell], page 59.

bind [-c class] key [command [args]]

Bind a command to a key. See Section 14.1 [Bind], page 73.

bindkey [opts] [string [cmd args]]

Bind a string to a series of keystrokes. See Section 14.5 [Bindkey], page 75.

blanker Blank the screen. See Section 20.23 [Screen Saver], page 96.

blankerprg

Define a blanker program. See Section 20.23 [Screen Saver], page 96.

break [duration]

Send a break signal to the current window. See Section 20.2 [Break], page 91.

breaktype [tcsendbreak | TCSBRK | TIOCSBRK]

Specify how to generate breaks. See Section 20.2 [Break], page 91.

bufferfile [exchange-file]

Select a file for screen-exchange. See Section 12.4 [Screen Exchange], page 69.

bumpleft Swaps window with previous one on window list. See Section 20.3 [Bumpleft], page 92.

bumpright

Swaps window with previous one on window list. See Section 20.4 [Bumpright], page 92.

c1 [state]

Change c1 code processing. See Section 11.11 [Character Processing], page 62.

caption mode [string]

Change caption mode and string. See Chapter 9 [Regions], page 39.

chacl usernames permbits list

Synonym to aclchg. See Section 8.4 [Multiuser Session], page 34.

charset set

Change character set slot designation. See Section 11.11 [Character Processing], page 62.

chdir [directory]

Change the current directory for future windows. See Section 6.1 [Chdir], page 27. cjkwidth Treat ambiguous width characters as full/half width. See Section 11.11 [Char-

acter Processing], page 62.

clear Clear the window screen. See Section 11.5 [Clear], page 60.

colon Enter a screen command. See Section 4.3 [Colon], page 9.

collapse Reorders window on window list, removing number gaps between them. See Section 20.5 [Collapse], page 92.

command [-c class]

Simulate the screen escape key. See Section 14.3 [Command Character], page 74.

compacthist [state]

Selects compaction of trailing empty lines. See Section 12.1.2 [Scrollback], page 65.

console [state]

Grab or ungrab console output. See Section 10.2 [Console], page 47.

copy Enter copy mode. See Section 12.1 [Copy], page 65.

copy_reg [key]

Removed. Use paste instead. See Section 12.3 [Registers], page 69.

crlf state

Select line break behavior for copying. See Section 12.1.1 [Line Termination], page 65.

debug state

Suppress/allow debugging output. See Section 20.6 [Debug], page 92.

defautonuke state

Select default autonuke behavior. See Section 16.6 [Autonuke], page 83.

defbce state

Select background color erase. See Section 11.11 [Character Processing], page 62.

defbreaktype [tcsendbreak | TCSBRK | TIOCSBRK]

Specify the default for generating breaks. See Section 20.2 [Break], page 91.

defc1 state

Select default c1 processing behavior. See Section 11.11 [Character Processing], page 62.

defcharset [set]

Change defaul character set slot designation. See Section 11.11 [Character Processing], page 62.

defencoding enc

Select default window encoding. See Section 11.11 [Character Processing], page 62.

defescape xy

Set the default command and meta characters. See Section 14.3 [Command Character], page 74.

defflow fstate

Select default flow control behavior. See Section 15.2 [Flow], page 77.

defgr state

Select default GR processing behavior. See Section 11.11 [Character Processing], page 62.

defhstatus [status]

Select default window hardstatus line. See Section 10.8 [Hardstatus], page 49.

deflog state

Select default window logging behavior. See Section 18.2 [Log], page 87.

deflogin state

Select default utmp logging behavior. See Section 10.4 [Login], page 47.

defmode mode

Select default file mode for ptys. See Section 10.5 [Mode], page 47.

defmonitor state

Select default activity monitoring behavior. See Section 10.6 [Monitor], page 48.

defmousetrack on|off

Select the default mouse tracking behavior. See Section 10.9 [Mousetrack], page 49.

defnonblock state|numsecs

Select default nonblock mode. See Section 20.9 [Nonblock], page 93.

defobuflimit limit

Select default output buffer limit. See Section 16.7 [Obuflimit], page 83.

defscrollback num

Set default lines of scrollback. See Section 12.1.2 [Scrollback], page 65.

defshell command

Set the default program for new windows. See Section 6.4 [Shell], page 28.

defsilence state

Select default idle monitoring behavior. See Section 10.6 [Monitor], page 48.

defslowpaste msec

Select the default inter-character timeout when pasting. See Section 12.2 [Paste], page 68.

defutf8 state

Select default character encoding. See Section 11.11 [Character Processing], page 62.

defwrap state

Set default line-wrapping behavior. See Section 11.8 [Wrap], page 61.

defwritelock on|off|auto

Set default writelock behavior. See Section 8.4 [Multiuser Session], page 34.

zombie_timeout [seconds]

Try to reconnect dead windows after timeout. See Section 20.14 [Zombie], page 94.

detach [-h]

Disconnect screen from the terminal. See Section 8.1 [Detach], page 33.

digraph [preset [unicode-value]]

Enter a digraph sequence. See Section 11.3 [Digraph], page 55.

dinfo Display terminal information. See Section 11.6 [Info], page 60.

displays List currently active user interfaces. See Section 8.4.6 [Displays], page 35.

dumptermcap

Write the window’s termcap entry to a file. See Section 16.2 [Dump Termcap], page 80.

echo [-n] message

Display a message on startup. See Chapter 19 [Startup], page 89.

encoding enc [denc]

Set the encoding of a window. See Section 11.11 [Character Processing], page 62.

escape xy Set the command and meta characters. See Section 14.3 [Command Character], page 74.

eval command1 [command2 ...]

Parse and execute each argument. See Section 20.20 [Eval], page 96.

exec [[fdpat] command [args ...]]

Run a subprocess (filter). See Section 13.1 [Exec], page 71.

fit Change window size to current display size. See Section 11.10 [Window Size], page 62.

flow [fstate]

Set flow control behavior. See Section 15.2 [Flow], page 77.

focus [next|prev|up|down|left|right|top|bottom]

Move focus to next region. See Chapter 9 [Regions], page 39.

focusminsize

Force the current region to a certain size. See Section 9.8 [Focusminsize], page 41.

gr [state]

Change GR charset processing. See Section 11.11 [Character Processing], page 62.

group [grouptitle]

Change or show the group the current window belongs to. See Section 6.7 [Window Groups], page 30.

hardcopy [-h] [file]

Write out the contents of the current window. See Section 18.1 [Hardcopy], page 87.

hardcopy_append state

Append to hardcopy files. See Section 18.1 [Hardcopy], page 87.

hardcopydir directory

Place, where to dump hardcopy files. See Section 18.1 [Hardcopy], page 87.

hardstatus [state]

Use the hardware status line. See Section 17.2 [Hardware Status Line], page 85.

height [lines [cols]]

Set display height. See Section 11.10 [Window Size], page 62.

help [-c class]

Display current key bindings. See Section 14.4 [Help], page 74.

history Find previous command beginning . . .. See Section 12.5 [History], page 70.

hstatus status

Change the window’s hardstatus line. See Section 10.8 [Hardstatus], page 49.

idle [timeout [cmd args]]

Define a screen saver command. See Section 20.23 [Screen Saver], page 96.

ignorecase [on|off]

Ignore character case in searches. See Section 12.1.7 [Searching], page 67.

info Display window settings. See Section 11.6 [Info], page 60.

ins_reg [key]

Removed, use paste instead. See Section 12.3 [Registers], page 69.

kill Destroy the current window. See Section 10.3 [Kill], page 47. lastmsg Redisplay the last message. See Section 17.3 [Last Message], page 85.

layout new [title]

Create a layout. See Section 9.9 [Layout], page 42.

layout remove [n|title]

Delete a layout. See Section 9.9 [Layout], page 42.

layout next

Select the next layout. See Section 9.9 [Layout], page 42.

layout prev

Select the previous layout. See Section 9.9 [Layout], page 42.

layout select [n|title]

Jump to a layout. See Section 9.9 [Layout], page 42.

layout show

List the available layouts. See Section 9.9 [Layout], page 42.

layout title [title]

Show or set the title of a layout. See Section 9.9 [Layout], page 42.

layout number [n]

Show or set the number of a layout. See Section 9.9 [Layout], page 42.

layout attach [title|:last]

Show or set which layout to reattach to. See Section 9.9 [Layout], page 42.

layout save [n|title]

Remember the organization of a layout. See Section 9.9 [Layout], page 42.

layout autosave [on|off]

Show or set the status of layout saving. See Section 9.9 [Layout], page 42.

layout dump [filename]

Save the layout arrangement to a file. See Section 9.9 [Layout], page 42.

license Display licensing information. See Chapter 19 [Startup], page 89.

lockscreen

Lock the controlling terminal. See Section 8.3 [Lock], page 34.

log [state]

Log all output in the current window. See Section 18.2 [Log], page 87.

logfile filename

Place where to collect logfiles. See Section 18.2 [Log], page 87.

login [state]

Log the window in /etc/utmp. See Section 10.4 [Login], page 47.

logtstamp [state]

Configure logfile time-stamps. See Section 18.2 [Log], page 87.

mapdefault

Use only the default mapping table for the next keystroke. See Section 14.7 [Bindkey Control], page 75.

mapnotnext

Don’t try to do keymapping on the next keystroke. See Section 14.7 [Bindkey Control], page 75.

maptimeout n

Set the inter-character timeout used for keymapping. See Section 14.7 [Bindkey Control], page 75.

markkeys string

Rebind keys in copy mode. See Section 12.1.3 [Copy Mode Keys], page 65.

maxwin n Set the maximum window number. See Section 20.21 [Maxwin], page 96.

meta Insert the command character. See Section 14.3 [Command Character], page 74.

monitor [state]

Monitor activity in window. See Section 10.6 [Monitor], page 48.

mousetrack [on|off]

Enable selecting split regions with mouse clicks. See Section 10.9 [Mousetrack], page 49.

msgminwait sec

Set minimum message wait. See Section 17.4 [Message Wait], page 86.

msgwait sec

Set default message wait. See Section 17.4 [Message Wait], page 86.

multiuser state

Go into single or multi user mode. See Section 8.4 [Multiuser Session], page 34.

nethack state

Use nethack-like error messages. See Section 20.8 [Nethack], page 93.

next Switch to the next window. See Chapter 7 [Selecting], page 31.

nonblock [state|numsecs]

Disable flow control to the current display. See Section 20.9 [Nonblock], page 93.|numsecs]

number [n]

Change/display the current window’s number. See Section 20.10 [Number], page 93.

obuflimit [limit]

Select output buffer limit. See Section 16.7 [Obuflimit], page 83.

only Kill all other regions. See Chapter 9 [Regions], page 39.

other Switch to the window you were in last. See Chapter 7 [Selecting], page 31.

partial state

Set window to partial refresh. See Section 11.7 [Redisplay], page 61.

password [crypted_pw]

Set reattach password. See Section 8.1 [Detach], page 33.

paste [src_regs [dest_reg]]

Paste contents of paste buffer or registers somewhere. See Section 12.2 [Paste], page 68.

pastefont [state]

Include font information in the paste buffer. See Section 12.2 [Paste], page 68.

pow_break

Close and Reopen the window’s terminal. See Section 20.2 [Break], page 91.

pow_detach

Detach and hang up. See Section 8.2 [Power Detach], page 33.

pow_detach_msg [message]

Set message displayed on pow_detach. See Section 8.2 [Power Detach], page 33.

prev Switch to the previous window. See Chapter 7 [Selecting], page 31.

printcmd [cmd]

Set a command for VT100 printer port emulation. See Section 20.15 [Printcmd], page 94.

process [key]

Treat a register as input to screen. See Section 12.3 [Registers], page 69.

quit Kill all windows and exit. See Section 8.7 [Quit], page 38.

readbuf [-e encoding] [filename]

Read the paste buffer from the screen-exchange file. See Section 12.4 [Screen Exchange], page 69.

readreg [-e encoding] [reg [file]]

Load a register from paste buffer or file. See Section 12.3 [Registers], page 69.

redisplay

Redisplay the current window. See Section 11.7 [Redisplay], page 61.

register [-e encoding] key string

Store a string to a register. See Section 12.3 [Registers], page 69.

remove Kill current region. See Chapter 9 [Regions], page 39.

removebuf

Delete the screen-exchange file. See Section 12.4 [Screen Exchange], page 69.

rendition bell | monitor | silence | so attr [color]

Change text attributes in caption for flagged windows. See Section 20.16 [Rendition], page 95.

reset Reset the terminal settings for the window. See Section 11.9 [Reset], page 62.

resize [(+/-)lines]

Grow or shrink a region. See Section 9.5 [Resize], page 40.

screen [opts] [n] [cmd [args] | //group]

Create a new window. See Section 6.2 [Screen Command], page 27.

scrollback num

Set size of scrollback buffer. See Section 12.1.2 [Scrollback], page 65.

select [n|-|.]

Switch to a specified window. See Chapter 7 [Selecting], page 31.

sessionname [name]

Name this session. See Section 8.5 [Session Name], page 37.

setenv [var [string]]

Set an environment variable for new windows. See Section 6.3 [Setenv], page 28.

setsid state

Controll process group creation for windows. See Section 20.19 [Setsid], page 95.

shell command

Set the default program for new windows. See Section 6.4 [Shell], page 28.

shelltitle title

Set the default name for new windows. See Section 6.4 [Shell], page 28.

silence [state|seconds]

Monitor a window for inactivity. See Section 10.6 [Monitor], page 48.

silencewait seconds

Default timeout to trigger an inactivity notify. See Section 10.6 [Monitor], page 48.

sleep num Pause during startup. See Chapter 19 [Startup], page 89.

slowpaste msec

Slow down pasting in windows. See Section 12.2 [Paste], page 68.

source file

Run commands from a file. See Section 4.2 [Source], page 9.

sorendition [attr [color]]

Deprecated. Use rendition so instead. See Section 20.16 [Rendition], page 95.

split Split region into two parts. See Chapter 9 [Regions], page 39.

startup_message state

Display copyright notice on startup. See Chapter 19 [Startup], page 89.

stuff [string]

Stuff a string in the input buffer of a window. See Section 12.2 [Paste], page 68.

su [username [password [password2]]]

Identify a user. See Section 8.4 [Multiuser Session], page 34.

suspend Put session in background. See Section 8.6 [Suspend], page 38.

term term Set $TERM for new windows. See Section 6.5 [Term], page 28.

termcap term terminal-tweaks [window-tweaks]

Tweak termcap entries for best performance. See Section 16.3 [Termcap Syntax], page 80.

terminfo term terminal-tweaks [window-tweaks]

Ditto, for terminfo systems. See Section 16.3 [Termcap Syntax], page 80.

termcapinfo term terminal-tweaks [window-tweaks]

Ditto, for both systems. See Section 16.3 [Termcap Syntax], page 80.

time [string]

Display time and load average. See Section 20.11 [Time], page 93.

title [windowtitle]

Set the name of the current window. See Section 10.1.1 [Title Command], page 45.

umask [users]+/-bits ...

Synonym to aclumask. See Section 8.4.7 [Umask], page 36.

unbindall

Unset all keybindings. See Section 14.1 [Bind], page 73.

unsetenv var

Unset environment variable for new windows. See Section 6.3 [Setenv], page 28.

utf8 [state [dstate]]

Select character encoding of the current window. See Section 11.11 [Character Processing], page 62.

vbell [state]

Use visual bell. See Section 11.4 [Bell], page 59.

vbell_msg [message]

Set vbell message. See Section 11.4 [Bell], page 59.

vbellwait sec

Set delay for vbell message. See Section 11.4 [Bell], page 59.

version Display screen version. See Section 20.13 [Version], page 94.

wall message

Write a message to all displays. See Section 8.4 [Multiuser Session], page 34.

width [cols [lines]]

Set the width of the window. See Section 11.10 [Window Size], page 62.

windowlist [[-b] [-m] [-g]] | string [string] | title [title]

Present a list of all windows for selection. See Section 7.4 [Windowlist], page 31.

windows List active windows. See Section 10.7 [Windows], page 49.

wrap [ on | off ]

Control line-wrap behavior. See Section 11.8 [Wrap], page 61.

writebuf [-e encoding] [filename]

Write paste buffer to screen-exchange file. See Section 12.4 [Screen Exchange], page 69.

writelock on|off|auto

Grant exclusive write permission. See Section 8.4 [Multiuser Session], page 34.

xoff Send an XOFF character. See Section 15.3 [XON/XOFF], page 78. xon Send an XON character. See Section 15.3 [XON/XOFF], page 78.

zmodem [off|auto|catch|pass]

Define how screen treats zmodem requests. See Section 20.24 [Zmodem], page 97.

zombie [keys [onerror] ]

Keep dead windows. See Section 20.14 [Zombie], page 94.

6 New Window

This section describes the commands for creating a new window for running programs. When a new window is created, the first available number is assigned to it. The number of windows is limited at compile-time by the MAXWIN configuration parameter (which defaults to 40).

6.1 Chdir

[Command] chdir [directory]

(none) Change the current directory of screen to the specified directory or, if called without an argument, to your home directory (the value of the environment variable $HOME). All windows that are created by means of the screen command from within .screenrc or by means of C-a : screen ... or C-a c use this as their default directory. Without a chdir command, this would be the directory from which screen was invoked. Hardcopy and log files are always written to the window’s default directory, not the current directory of the process running in the window. You can use this command multiple times in your .screenrc to start various windows in different default directories, but the last chdir value will affect all the windows you create interactively.

6.2 Screen Command

[Command] screen [opts] [n] [cmd [args] | //group]

(C-a c, C-a C-c)

Establish a new window. The flow-control options (‘-f’, ‘-fn’ and ‘-fa’), title option (‘-t’), login options (‘-l’ and ‘-ln’) , terminal type option (‘-T term’), the all-capability-flag (‘-a’) and scrollback option (‘-h num’) may be specified with each command. The option (‘-M’) turns monitoring on for this window. The option (‘-L’) turns output logging on for this window. If an optional number n in the range 0…MAXWIN-1 is given, the window number n is assigned to the newly created window (or, if this number is already in-use, the next available number). If a command is specified after screen, this command (with the given arguments) is started in the window; otherwise, a shell is created. If ‘//group’ is supplied, a container-type window is created in which other windows may be created inside it. See Section 6.7 [Window Groups], page 30. Screen has built in some functionality of ‘cu’ and ‘telnet’. See Section 6.6 [Window Types], page 29.

Thus, if your .screenrc contains the lines

# example for .screenrc:
screen 1
screen -fn -t foobar 2 -L telnet foobar

screen creates a shell window (in window #1) and a window with a TELNET connection to the machine foobar (with no flow-control using the title ‘foobar’ in window #2) and will write a logfile ‘screenlog.2’ of the telnet session. If you do not include any screen commands in your .screenrc file, then screen defaults to creating a single shell window, number zero. When the initialization is completed, screen switches to the last window specified in your .screenrc file or, if none, it opens default window #0.

6.3 Setenv

[Command] setenv var string

(none) Set the environment variable var to value string. If only var is specified, the user will be prompted to enter a value. If no parameters are specified, the user will be prompted for both variable and value. The environment is inherited by all subsequently forked shells.

[Command] unsetenv var

(none) Unset an environment variable.

6.4 Shell

[Command] shell command

[Command] defshell command

(none) Set the command to be used to create a new shell. This overrides the value of the environment variable $SHELL. This is useful if you’d like to run a tty-enhancer which is expecting to execute the program specified in $SHELL. If the command begins with a ‘-’ character, the shell will be started as a login-shell. Typical shells do only minimal initialization when not started as a login-shell. E.g. Bash will not read your ~/.bashrc unless it is a login-shell. defshell is currently a synonym to the shell .screenrc command.

[Command] shelltitle title

(none) Set the title for all shells created during startup or by the C-a C-c command. See Section 10.1 [Naming Windows], page 45, for details about what titles are.

6.5 Term

[Command] term term

(none) In each window screen opens, it sets the $TERM variable to screen by default, unless no description for screen is installed in the local termcap or terminfo data base. In that case it pretends that the terminal emulator is ‘vt100’. This won’t do much harm, as screen is VT100/ANSI compatible. The use of the term command is discouraged for non-default purpose. That is, one may want to specify special $TERM settings (e.g. vt100) for the next screen rlogin othermachine command. Use the command screen -T vt100 rlogin othermachine rather than setting and resetting the default.

6.6 Window Types

Screen provides three different window types. New windows are created with screen’s ‘screen’ command (see Section 6.2 [Screen Command], page 27). The first parameter to the ‘screen’ command defines which type of window is created. The different window types are all special cases of the normal type. They have been added in order to allow screen to be used efficiently as a console with 100 or more windows.

• The normal window contains a shell (default, if no parameter is given) or any other system command that could be executed from a shell. (e.g. ‘slogin’, etc...).

• If a tty (character special device) name (e.g. ‘/dev/ttya’) is specified as the first parameter, then the window is directly connected to this device. This window type is similar to ‘screen cu -l /dev/ttya’. Read and write access is required on the device node, an exclusive open is attempted on the node to mark the connection line as busy. An optional parameter is allowed consisting of a comma separated list of flags in the notation used by ‘stty(1)’:

<baud_rate>

Usually 300, 1200, 9600 or 19200. This affects transmission as well as receive speed.

cs8 or cs7

Specify the transmission of eight (or seven) bits per byte.

ixon or -ixon

Enables (or disables) software flow-control (CTRL-S/CTRL-Q) for sending data.

ixoff or -ixoff

Enables (or disables) software flow-control for receiving data.

istrip or -istrip

Clear (or keep) the eight bit in each received byte.

You may want to specify as many of these options as applicable. Unspecified options cause the terminal driver to make up the parameter values of the connection. These values are system-dependent and may be in defaults or values saved from a previous connection. For tty windows, the info command shows some of the modem control lines in the status line. These may include ‘RTS’, ‘CTS’, ‘DTR’, ‘CD’ and more. This depends rather on on the available ioctl()’s and system header files than on the physical capabilities of the serial board. The name of a logical low (inactive) signal is preceded by an exclamation mark (‘!’), otherwise the signal is logical high (active). Unsupported but shown signals are usually shown low. When the CLOCAL status bit is true, the whole set of modem signals is placed inside curly braces (‘{’ and ‘}’). When the CRTSCTS or TIOCSOFTCAR bit is true, the signals ‘CTS’ or ‘CD’ are shown in parenthesis, respectively. For tty windows, the command break causes the Data transmission line (TxD) to go low for a specified period of time. This is expected to be interpreted as break signal on the other side. No data is sent and no modem control line is changed when a break is issued.

• If the first parameter is //telnet, the second parameter is expected to be a host name, and an optional third parameter may specify a TCP port number (default decimal 23). Screen will connect to a server listening on the remote host and use the telnet protocol to communicate with that server. For telnet windows, the command info shows details about the connection in square brackets (‘[’ and ‘]’) at the end of the status line.

b BINARY. The connection is in binary mode.

e ECHO. Local echo is disabled.

c SGA. The connection is in ‘character mode’ (default: ‘line mode’).

t TTYPE. The terminal type has been requested by the remote host. Screen sends the name screen unless instructed otherwise (see also the command

          ‘term’).

w NAWS. The remote site is notified about window size changes.

f LFLOW. The remote host will send flow control information. (Ignored at the moment.)

Additional flags for debugging are ‘x’, ‘t’ and ‘n’ (XDISPLOC, TSPEED and NEWENV). For telnet windows, the command break sends the telnet code IAC BREAK (decimal 243) to the remote host.

6.7 Window Groups

Screen provides a method for grouping windows together. Windows can be organized in a hierarchical fashion, resembling a tree structure. New screens are created using the screen command while new groups are created using screen //group. See Section 6.2 [Screen Command], page 27.

Once a new group is created, it will act as a container for windows and even other groups. When a group is selected, you will see the output of the windowlist command, allowing you to select a window inside. If there are no windows inside a group, use the screen command to create one. Once inside a group, using the commands next and prev will switch between windows only in that group. Using the windowlist command will give you the opportunity to leave the group you are in. See Section 7.4 [Windowlist], page 31.

[Command] group [grouptitle]

(none) Change or show the group the current window belongs to. Windows can be moved around between different groups by specifying the name of the destination group. Without specifying a group, the title of the current group is displayed.

Using groups in combination with layouts will help create a multi-desktop experience. One group can be assigned for each layout made. Windows can be made, split, and organized within each group as desired. Afterwhich, switching between groups can be as easy as switching layouts.

7 Selecting a Window

This section describes the commands for switching between windows in an screen session. The windows are numbered from 0 to 9, and are created in that order by default (see Chapter 6 [New Window], page 27).

7.1 Moving Back and Forth

[Command] next

(C-a SPC, C-a n, C-a C-n)

Switch to the next window. This command can be used repeatedly to cycle through the list of windows. (On some terminals, C-SPC generates a NUL character, so you must release the control key before pressing space.)

[Command] prev

(C-a p, C-a C-p, C-a C-h, C-a Backspace)

Switch to the previous window (the opposite of C-a n).

7.2 Other Window

[Command] other

(C-a C-a)

Switch to the last window displayed. Note that this command defaults to the command character typed twice, unless overridden. For instance, if you use the option ‘-e]x’, this command becomes ]] (see Section 14.3 [Command Character], page 74).

7.3 Select

[Command] select [n |-|.]

(C-a n, C-a ')

Switch to the window with the number n. If no window number is specified, you get prompted for an identifier. This can be a window name (title) or a number. When a new window is established, the lowest available number is assigned to this window. Thus, the first window can be activated by select 0; there can be no more than 10 windows present simultaneously (unless screen is compiled with a higher MAXWIN setting). There are two special arguments, select - switches to the internal blank window and select . switches to the current window. The latter is useful if used with screen’s -X option.

7.4 Windowlist

[Command] windowlist [-b] [-m] [-g]

[Command] windowlist string [string]

[Command] windowlist title [title]

(C-a ")

Display all windows in a table for visual window selection. If screen was in a window group, screen will back out of the group and then display the windows in that group.

If the -b option is given, screen will switch to the blank window before presenting the list, so that the current window is also selectable. The -m option changes the order of the windows, instead of sorting by window numbers screen uses its internal most-recently-used list. The -g option will show the windows inside any groups in that level and downwards. The following keys are used to navigate in windowlist: k, C-p, or up Move up one line. j, C-n, or down Move down one line. C-g or escape Exit windowlist. C-a or home Move to the first line. C-e or end Move to the last line. C-u or C-d Move one half page up or down. C-b or C-f Move one full page up or down. 0..9 Using the number keys, move to the selected line. mouseclick Move to the selected line. Available when mousetrack is set to on. / Search. n Repeat search in the forward direction. N Repeat search in the backward direction. m Toggle MRU. g Toggle group nesting. a All window view. C-h or backspace Back out the group. , Switch numbers with the previous window. . Switch numbers with the next window. K Kill that window. space or enter Select that window. The table format can be changed with the string and title option, the title is displayed as table heading, while the lines are made by using the string setting. The default setting is ‘Num Name%=Flags’ for the title and ‘%3n %t%=%f’ for the lines. See the string escapes chapter (see Chapter 21 [String Escapes], page 99) for more codes (e.g. color settings). Windowlist needs a region size of at least 10 characters wide and 6 characters high in order to display.

8 Session Management Commands

Perhaps the most useful feature of screen is the way it allows the user to move a session between terminals, by detaching and reattaching. This also makes life easier for modem users who have to deal with unexpected loss of carrier.

8.1 Detach

[Command] autodetach state

(none) Sets whether screen will automatically detach upon hangup, which saves all your running programs until they are resumed with a screen -r command. When turned off, a hangup signal will terminate screen and all the processes it contains. Autodetach is on by default.

[Command] detach

(C-a d, C-a C-d)

Detach the screen session (disconnect it from the terminal and put it into the background). A detached screen can be resumed by invoking screen with the -r option (see Chapter 3 [Invoking Screen], page 5). The -h option tells screen to immediately close the connection to the terminal (‘hangup’).

[Command] password [crypted pw]

(none) Present a crypted password in your .screenrc file and screen will ask for it, whenever someone attempts to resume a detached session. This is useful, if you have privileged programs running under screen and you want to protect your session from reattach attempts by users that managed to assume your uid. (I.e. any superuser.) If no crypted password is specified, screen prompts twice a password and places its encryption in the paste buffer. Default is ‘none’, which disables password checking.

8.2 Power Detach

[Command] pow_detach
(C-a D D)

Mainly the same as detach, but also sends a HANGUP signal to the parent process

of screen.

Caution: This will result in a logout if screen was started from your login-shell.

[Command] pow_detach_msg [message]

(none) The message specified here is output whenever a power detach is performed. It may be used as a replacement for a logout message or to reset baud rate, etc. Without a parameter, the current message is shown.

8.3 Lock

[Command] lockscreen
(C-a x, C-a C-x)

Call a screenlock program (/local/bin/lck or /usr/bin/lock or a builtin, if no other is available). Screen does not accept any command keys until this program terminates. Meanwhile processes in the windows may continue, as the windows are in the detached state. The screenlock program may be changed through the environment variable $LOCKPRG (which must be set in the shell from which screen is started) and is executed with the user’s uid and gid. Warning: When you leave other shells unlocked and have no password set on screen, the lock is void: One could easily re-attach from an unlocked shell. This feature should rather be called lockterminal.

8.4 Multiuser Session

These commands allow other users to gain access to one single screen session. When attaching to a multiuser screen the sessionname is specified as username/sessionname to the -S command line option. Screen must be compiled with multiuser support to enable features described here.

8.4.1 Multiuser

[Command] multiuser state

(none) Switch between single-user and multi-user mode. Standard screen operation is singleuser. In multi-user mode the commands acladd, aclchg and acldel can be used to enable (and disable) other users accessing this screen.

8.4.2 Acladd

[Command] acladd usernames
[Command] addacl usernames

(none) Enable users to fully access this screen session. Usernames can be one user or a comma separated list of users. This command enables to attach to the screen session and performs the equivalent of aclchg usernames +rwx "#?". To add a user with restricted access, use the aclchg command below. Addacl is a synonym to acladd. Multi-user mode only.

8.4.3 Aclchg

[Command] aclchg usernames permbits list
[Command] chacl usernames permbits list

(none) Change permissions for a comma separated list of users. Permission bits are represented as ‘r’, ‘w’ and ‘x’. Prefixing ‘+’ grants the permission, ‘-’ removes it. The third parameter is a comma separated list of commands or windows (specified either by number or title). The special list ‘#’ refers to all windows, ‘?’ to all commands. If usernames consists of a single ‘*’, all known users are affected. A command can be executed when the user has the ‘x’ bit for it. The user can type input to a window when he has its ‘w’ bit set and no other user obtains a writelock for this window. Other bits are currently ignored. To withdraw the writelock from another user in e.g. window 2: ‘aclchg username -w+w 2’. To allow read-only access to the session: ‘aclchg username -w "#"’. As soon as a user’s name is known to screen, he can attach to the session and (per default) has full permissions for all command and windows. Execution permission for the acl commands, at and others should also be removed or the user may be able to regain write permission. Chacl is a synonym to aclchg. Multi-user mode only.

8.4.4 Acldel

[Command] acldel username

(none) Remove a user from screen’s access control list. If currently attached, all the user’s displays are detached from the session. He cannot attach again. Multi-user mode only.

8.4.5 Aclgrp

[Command] aclgrp username [groupname]

(none) Creates groups of users that share common access rights. The name of the group is the username of the group leader. Each member of the group inherits the permissions that are granted to the group leader. That means, if a user fails an access check, another check is made for the group leader. A user is removed from all groups the special value ‘none’ is used for groupname. If the second parameter is omitted all groups the user is in are listed.

8.4.6 Displays

[Command] displays

(C-a *)

Shows a tabular listing of all currently connected user front-ends (displays). This is most useful for multiuser sessions. The following keys can be used in displays list: k, C-p, or up Move up one line. j, C-n, or down Move down one line. C-a or home Move to the first line. C-e or end Move to the last line. C-u or C-d Move one half page up or down. C-b or C-f Move one full page up or down. mouseclick Move to the selected line. Available when mousetrack is set to on. space Refresh the list. d Detach the selected display.

D Power detach the selected display.

C-g, enter, or escape Exit the list.

The following is an example of what displays could look like:

     xterm 80x42   jnweiger@/dev/ttyp4      0(m11)    &rWx
     facit 80x24   mlschroe@/dev/ttyhf nb  11(tcsh)    rwx
     xterm 80x42   jnhollma@/dev/ttyp5      0(m11)    &R.x
      (A)   (B)       (C)     (D)     (E)  (F)(G)    (H)(I)

The legend is as follows: (A) The terminal type known by screen for this display. (B) Displays geometry as width x height. (C) Username who is logged in at the display. (D) Device name of the display or the attached device (E) Display is in blocking or nonblocking mode. The available modes are "nb", "NB", "Z<", "Z>", and "BL". (F) Number of the window (G) Name/title of window (H) Whether the window is shared (I) Window permissions. Made up of three characters:

(1st character) ‘-’ : no read ‘r’ : read ‘R’ : read only due to foreign wlock

(2nd character) ‘-’ : no write ‘.’ : write suppressed by foreign wlock ‘w’ : write ‘W’ : own wlock

(3rd character)

‘-’ : no execute ‘x’ : execute

Displays needs a region size of at least 10 characters wide and 5 characters high in order to display.

8.4.7 aclumask

[Command] aclumask [users]+/-bits ...

[Command] umask [users]+/-bits ...

(none) This specifies the access other users have to windows that will be created by the caller of the command. Users may be no, one or a comma separated list of known usernames. If no users are specified, a list of all currently known users is assumed. Bits is any combination of access control bits allowed defined with the aclchg command. The special username ‘?’ predefines the access that not yet known users will be granted to any window initially. The special username ‘??’ predefines the access that not yet known users are granted to any command. Rights of the special username nobody cannot be changed (see the su command). Umask is a synonym to aclumask.

8.4.8 Wall

[Command] wall message

(none) Write a message to all displays. The message will appear in the terminal’s status line.

8.4.9 Writelock

[Command] writelock on|off|auto

(none) In addition to access control lists, not all users may be able to write to the same window at once. Per default, writelock is in ‘auto’ mode and grants exclusive input permission to the user who is the first to switch to the particular window. When he leaves the window, other users may obtain the writelock (automatically). The writelock of the current window is disabled by the command writelock off. If the user issues the command writelock on he keeps the exclusive write permission while switching to other windows.

[Command] defwritelock on|off|auto

(none) Sets the default writelock behavior for new windows. Initially all windows will be created with no writelocks.

8.4.10 Su

[Command] su [username [password [password2]]]

(none) Substitute the user of a display. The command prompts for all parameters that are omitted. If passwords are specified as parameters, they have to be specified uncrypted. The first password is matched against the systems passwd database, the second password is matched against the screen password as set with the commands acladd or password. Su may be useful for the screen administrator to test multiuser setups. When the identification fails, the user has access to the commands available

for user ‘nobody’. These are detach, license, version, help and displays.

8.5 Session Name

[Command] sessionname [name]

(none) Rename the current session. Note that for screen -list the name shows up with the process-id prepended. If the argument name is omitted, the name of this session is displayed. Caution: The $STY environment variable will still reflect the old name in pre-existing shells. This may result in confusion. Use of this command is generally discouraged. Use the -S command-line option if you want to name a new session.The default is constructed from the tty and host names.

8.6 Suspend

[Command] suspend

(C-a z, C-a C-z)

Suspend screen. The windows are in the detached state while screen is suspended. This feature relies on the parent shell being able to do job control.

8.7 Quit

[Command] quit

(C-a \)

Kill all windows and terminate screen. (see Chapter 14 [Key Binding], page 73).

9 Regions

Screen has the ability to display more than one window on the user’s display. This is done by splitting the screen in regions, which can contain different windows.

9.1 Split

[Command] split [-v]

(C-a S, C-a |)

Split the current region into two new ones. All regions on the display are resized to make room for the new region. The blank window is displayed in the new region. The default is to create a horizontal split, putting the new regions on the top and bottom of each other. Using ‘-v’ will create a vertical split, causing the new regions to appear side by side of each other. Use the remove or the only command to delete regions. Use focus to toggle between regions. When a region is split opposite of how it was previously split (that is, vertical then horizontal or horizontal then vertical), a new layer is created. The layer is used to group together the regions that are split the same. Normally, as a user, you should not see nor have to worry about layers, but they will affect how some commands (focus and resize) behave. With this current implementation of screen, scrolling data will appear much slower in a vertically split region than one that is not. This should be taken into consideration if you need to use system commands such as cat or tail -f.

9.2 Focus

[Command] focus [ next|prev|up|down|left|right|top|bottom ]
(C-a Tab)

Move the input focus to the next region. This is done in a cyclic way so that the top left region is selected after the bottom right one. If no option is given it defaults to next. The next region to be selected is determined by how the regions are layered. Normally, the next region in the same layer would be selected. However, if that next region contains one or more layers, the first region in the highest layer is selected first. If you are at the last region of the current layer, next will move the focus to the next region in the lower layer (if there is a lower layer). Prev cycles in the opposite order. See Section 9.1 [Split], page 39, for more information about layers. The rest of the options (up, down, left, right, top, and bottom) are more indifferent to layers. The option up will move the focus upward to the region that is touching the upper left corner of the current region. Down will move downward to the region that is touching the lower left corner of the current region. The option left will move the focus leftward to the region that is touching the upper left corner of the current region, while right will move rightward to the region that is touching the upper right corner of the current region. Moving left from a left most region or moving right from a right most region will result in no action. The option top will move the focus to the very first region in the upper list corner of the screen, and bottom will move to the region in the bottom right corner of the screen. Moving up from a top most region or moving down from a bottom most region will result in no action. Useful bindings are (h, j, k, and l as in vi):

     bind h focus left
     bind j focus down
     bind k focus up
     bind l focus right
     bind t focus top
     bind b focus bottom

Note that ‘k’ is traditionally bound to the kill command.

9.3 Only

[Command] only

(C-a Q)

Kill all regions but the current one.

9.4 Remove

[Command] remove

(C-a X)

Kill the current region. This is a no-op if there is only one region.

9.5 Resize

[Command] resize [-h|-v|-b|-l|-p] [ [+|-]n[%] | = | max | min | _| 0 ]

(none) Resize the current region. The space will be removed from or added to the surrounding regions depending on the order of the splits. The available options for resizing are ‘-h’(horizontal), ‘-v’(vertical), ‘-b’(both), ‘-l’(local to layer), and ‘-p’(perpendicular). Horizontal resizes will add or remove width to a region, vertical will add or remove height, and both will add or remove size from both dimensions. Local and perpendicular are similar to horizontal and vertical, but they take in account of how a region was split. If a region’s last split was horizontal, a local resize will work like a vertical resize. If a region’s last split was vertical, a local resize will work like a horizontal resize. Perpendicular resizes work in opposite of local resizes. If no option is specified, local is the default. The amount of lines to add or remove can be expressed a couple of different ways. By specifying a number n by itself will resize the region by that absolute amount. You can specify a relative amount by prefixing a plus ‘+’ or minus ‘-’ to the amount, such as adding +n lines or removing -n lines. Resizing can also be expressed as an absolute or relative percentage by postfixing a percent sign ‘%’. Using zero ‘0’ is a synonym for min and using an underscore ‘_’ is a synonym for max. Some examples are:

     resize +N       increase current region by N

     resize -N       decrease current region by N
     resize N        set current region to N
     resize 20%      set current region to 20% of original size
     resize +20%     increase current region by 20%
     resize -b =     make all windows equally
     resize max      maximize current region
     resize min      minimize current region

Without any arguments, screen will prompt for how you would like to resize the current region.

See focusminsize if you want to restrict the minimun size a region can have.

9.6 Caption

[Command] caption always|splitonly [string]

[Command] caption string [string]

(none) This command controls the display of the window captions. Normally a caption is only used if more than one window is shown on the display (split screen mode). But if the type is set to always, screen shows a caption even if only one window is displayed. The default is ‘splitonly’.

The second form changes the text used for the caption. You can use all string escapes (see Chapter 21 [String Escapes], page 99). Screen uses a default of ‘%3n %t’.

You can mix both forms by providing the string as an additional argument.

9.7 Fit

[Command] fit

(C-a F)

Change the window size to the size of the current region. This command is needed because screen doesn’t adapt the window size automatically if the window is displayed more than once.

9.8 Focusminsize

[Command] focusminsize [ (width|max|_) (height|max|_) ]

(none) This forces any currently selected region to be automatically resized at least a certain width and height. All other surrounding regions will be resized in order to accommodate. This constraint follows every time the focus command is used. The resize command can be used to increase either dimension of a region, but never below what is set with focusminsize. The underscore ‘_’ is a synonym for max. Setting a width and height of 0 0 (zero zero) will undo any constraints and allow for manual resizing. Without any parameters, the minimum width and height is shown.

9.9 Layout

Using regions, and perhaps a large enough terminal, you can give screen more of a desktop feel. By being able to split regions horizontally or vertically, you can take advantage of the lesser used spaces of your terminal. The catch to these splits has been that they’re not kept between screen detachments and reattachments.

Layouts will help organize your regions. You can create one layout of four horizontal regions and then create a separate layout of regions in a two by two array. The regions could contain the same windows, but they don’t have to. You can easily switch between layouts and keep them between detachments and reattachments.

Note that there are several subcommands to layout.

[Command] layout new [title]

(none) Create a new layout. The screen will change to one whole region and be switched to the blank window. From here, you build the regions and the windows they show as you desire. The new layout will be numbered with the smallest available integer, starting with zero. You can optionally give a title to your new layout. Otherwise, it will have a default title of layout. You can always change the title later by using the

command layout title.

[Command] layout remove [n|title]

(none) Remove, or in other words, delete the specified layout. Either the number or the title can be specified. Without either specification, screen will remove the current layout. Removing a layout does not affect your set windows or regions.

[Command] layout next

(none) Switch to the next layout available

[Command] layout prev

(none) Switch to the previous layout available

[Command] layout select [n|title]

(none) Select the desired layout. Either the number or the title can be specified. Without either specification, screen will prompt and ask which screen is desired. To see which layouts are available, use the layout show command.

[Command] layout show

(none) List on the message line the number(s) and title(s) of the available layout(s). The current layout is flagged.

[Command] layout title [title]

(none) Change or display the title of the current layout. A string given will be used to name the layout. Without any options, the current title and number is displayed on the message line.

[Command] layout number [n]

(none) Change or display the number of the current layout. An integer given will be used to number the layout. Without any options, the current number and title is displayed on the message line.

[Command] layout attach [title|:last]

(none) Change or display which layout to reattach back to. The default is :last, which tells screen to reattach back to the last used layout just before detachment. By supplying a title, You can instruct screen to reattach to a particular layout regardless which one was used at the time of detachment. Without any options, the layout to reattach to will be shown in the message line.

[Command] layout save [n|title]

(none) Remember the current arrangement of regions. When used, screen will remember the arrangement of vertically and horizontally split regions. This arrangement is restored when a screen session is reattached or switched back from a different layout. If the session ends or the screen process dies, the layout arrangements are lost. The layout dump command should help in this siutation. If a number or title is supplied, screen will remember the arrangement of that particular layout. Without any options, screen will remember the current layout. Saving your regions can be done automatically by using the layout autosave command.

[Command] layout autosave [on|off]

(none) Change or display the status of automatically saving layouts. The default is on, meaning when screen is detached or changed to a different layout, the arrangement of regions and windows will be remembered at the time of change and restored upon return. If autosave is set to off, that arrangement will only be restored to either to the last manual save, using layout save, or to when the layout was first created, to a single region with a single window. Without either an on or an off, the current status is displayed on the message line.

[Command] layout dump [filename]

(none) Write to a file the order of splits made in the current layout. This is useful to recreate the order of your regions used in your current layout. Only the current layout is recorded. While the order of the regions are recorded, the sizes of those regions and which windows correspond to which regions are not. If no filename is specified, the default is layout-dump, saved in the directory that the screen process was started in. If the file already exists, layout dump will append to that file. As an example:

     layout dump /home/user/.screenrc

will save or append the layout to the user’s .screenrc file.

10 Window Settings

These commands control the way screen treats individual windows in a session. See Chapter 11 [Virtual Terminal], page 51, for commands to control the terminal emulation itself.

10.1 Naming Windows (Titles)

You can customize each window’s name in the window display (viewed with the windows command (see Section 10.7 [Windows], page 49) by setting it with one of the title commands. Normally the name displayed is the actual command name of the program created in the window. However, it is sometimes useful to distinguish various programs of the same name or to change the name on-the-fly to reflect the current state of the window.

The default name for all shell windows can be set with the shelltitle command (see Section 6.4 [Shell], page 28). You can specify the name you want for a window with the ‘-t’ option to the screen command when the window is created (see Section 6.2 [Screen Command], page 27). To change the name after the window has been created you can use the title-string escape-sequence (ESC k name ESC \) and the title command (C-a A). The former can be output from an application to control the window’s name under software control, and the latter will prompt for a name when typed. You can also bind predefined names to keys with the title command to set things quickly without prompting.

10.1.1 Title Command

[Command] title [windowtitle]

     (C-a A)

Set the name of the current window to windowtitle. If no name is specified, screen prompts for one.

10.1.2 Dynamic Titles

screen has a shell-specific heuristic that is enabled by setting the window’s name to search|name and arranging to have a null title escape-sequence output as a part of your prompt. The search portion specifies an end-of-prompt search string, while the name portion specifies the default shell name for the window. If the name ends in a ‘:’ screen will add what it believes to be the current command running in the window to the end of the specified name (e.g. name:cmd). Otherwise the current command name supersedes the shell name while it is running.

Here’s how it works: you must modify your shell prompt to output a null title-escape-sequence (ESC k ESC \) as a part of your prompt. The last part of your prompt must be the same as the string you specified for the search portion of the title. Once this is set up, screen will use the title-escape-sequence to clear the previous command name and get ready for the next command. Then, when a newline is received from the shell, a search is made for the end of the prompt. If found, it will grab the first word after the matched string and use it as the command name. If the command name begins with ‘!’, ‘%’, or ‘^’, screen will use the first word on the following line (if found) in preference to the just-found name. This helps csh users get more accurate titles when using job control or history recall commands.

10.1.3 Setting up your prompt for shell titles

One thing to keep in mind when adding a null title-escape-sequence to your prompt is that some shells (like the csh) count all the non-control characters as part of the prompt’s length. If these invisible characters aren’t a multiple of 8 then backspacing over a tab will result in an incorrect display. One way to get around this is to use a prompt like this:

     set prompt='<ESC>[0000m<ESC>k<ESC>\% '

The escape-sequence ‘<ESC>[0000m’ not only normalizes the character attributes, but all the zeros round the length of the invisible characters up to 8.

Tcsh handles escape codes in the prompt more intelligently, so you can specify your prompt like this:

     set prompt="%{\ek\e\\%}\% "

Bash users will probably want to echo the escape sequence in the PROMPT COMMAND:

     PROMPT_COMMAND='printf "\033k\033\134"'

(I used ‘\134’ to output a ‘\’ because of a bug in v1.04).

10.1.4 Setting up shell titles in your .screenrc

Here are some .screenrc examples:

     screen -t top 2 nice top

Adding this line to your .screenrc would start a niced version of the top command in window 2 named ‘top’ rather than ‘nice’.

     shelltitle '> |csh'
     screen 1

This file would start a shell using the given shelltitle. The title specified is an auto-title that would expect the prompt and the typed command to look something like the following:

     /usr/joe/src/dir> trn

(it looks after the ’> ’ for the command name). The window status would show the name ‘trn’ while the command was running, and revert to ‘csh’ upon completion.

     bind R screen -t '% |root:' su

Having this command in your .screenrc would bind the key sequence C-a R to the su command and give it an auto-title name of ‘root:’. For this auto-title to work, the screen could look something like this:

     % !em
     emacs file.c

Here the user typed the csh history command !em which ran the previously entered emacs command. The window status would show ‘root:emacs’ during the execution of the command, and revert to simply ‘root:’ at its completion.

     bind o title
     bind E title ""
     bind u title (unknown)

The first binding doesn’t have any arguments, so it would prompt you for a title when you type C-a o. The second binding would clear an auto-titles current setting (C-a E). The third binding would set the current window’s title to ‘(unknown)’ (C-a u).

10.2 Console

[Command] console [state]

(none) Grabs or un-grabs the machines console output to a window. When the argument is omitted the current state is displayed. Note: Only the owner of /dev/console can grab the console output. This command is only available if the host supports the

ioctl TIOCCONS.

10.3 Kill

[Command] kill

(C-a k, C-a C-k)

Kill the current window. If there is an exec command running (see Section 13.1 [Exec], page 71) then it is killed. Otherwise the process (e.g. shell) running in the window receives a HANGUP condition, the window structure is removed and screen (your display) switches to another window. When the last window is destroyed, screen exits. After a kill screen switches to the previously displayed window. Caution: emacs users may find themselves killing their emacs session when trying to delete the current line. For this reason, it is probably wise to use a different command character (see Section 14.3 [Command Character], page 74) or rebind kill to another key sequence, such as C-a K (see Chapter 14 [Key Binding], page 73).

10.4 Login

[Command] deflogin state

(none) Same as the login command except that the default setting for new windows is changed. This defaults to ‘on’ unless otherwise specified at compile time (see Chapter 26 [Installation], page 111). Both commands are only present when screen has been compiled with utmp support.

[Command] login [state]

(C-a L)

Adds or removes the entry in /etc/utmp for the current window. This controls whether or not the window is logged in. In addition to this toggle, it is convenient to have “log in” and “log out” keys. For instance, bind I login on and bind O login off will map these keys to be C-a I and C-a O (see Chapter 14 [Key Binding], page 73).

10.5 Mode

[Command] defmode mode

(none) The mode of each newly allocated pseudo-tty is set to mode. mode is an octal number as used by chmod(1). Defaults to 0622 for windows which are logged in, 0600 for others (e.g. when -ln was specified for creation, see Section 6.2 [Screen Command], page 27).

10.6 Monitoring

[Command] activity message

(none) When any activity occurs in a background window that is being monitored, screen displays a notification in the message line. The notification message can be redefined by means of the activity command. Each occurrence of ‘%’ in message is replaced by the number of the window in which activity has occurred, and each occurrence of ‘^G’ is replaced by the definition for bell in your termcap (usually an audible bell). The default message is

     'Activity in window %n'

Note that monitoring is off for all windows by default, but can be altered by use of the monitor command (C-a M).

[Command] defmonitor state

(none) Same as the monitor command except that the default setting for new windows is changed. Initial setting is ‘off’.

[Command] monitor [state]

(C-a M)

Toggles monitoring of the current window. When monitoring is turned on and the affected window is switched into the background, the activity notification message will be displayed in the status line at the first sign of output, and the window will also be marked with an ‘@’ in the window-status display (see Section 10.7 [Windows], page 49). Monitoring defaults to ‘off’ for all windows.

[Command] silence [state|sec]
(C-a _)

Toggles silence monitoring of windows. When silence is turned on and an affected window is switched into the background, you will receive the silence notification message in the status line after a specified period of inactivity (silence). The default timeout can be changed with the silencewait command or by specifying a number of seconds instead of on or off. Silence is initially off for all windows.

[Command] defsilence state

(none) Same as the silence command except that the default setting for new windows is changed. Initial setting is ‘off’.

[Command] silencewait seconds

(none) Define the time that all windows monitored for silence should wait before displaying a message. Default is 30 seconds.

10.7 Windows

[Command] windows [ string ]

(C-a w, C-a C-w)

Uses the message line to display a list of all the windows. Each window is listed by number with the name of the program running in the window (or its title). The current window is marked with a ‘*’; the previous window is marked with a ‘-’; all the windows that are logged in are marked with a ‘$’ (see Section 10.4 [Login], page 47); a background window that has received a bell is marked with a ‘!’; a background window that is being monitored and has had activity occur is marked with an ‘@’ (see Section 10.6 [Monitor], page 48); a window which has output logging turned on is marked with ‘(L)’; windows occupied by other users are marked with ‘&’ or ‘&&’ if the window is shared by other users; windows in the zombie state are marked with ‘Z’. If this list is too long to fit on the terminal’s status line only the portion around the current window is displayed. You can customize the output format to any string you like including string escapes (see Chapter 21 [String Escapes], page 99). In this case, if the string parameter is passed, the maximum output size is unlimited (instead of 1024 bytes if no parameter is passed).

10.8 Hardstatus

Screen maintains a hardstatus line for every window. If a window gets selected, the display’s hardstatus will be updated to match the window’s hardstatus line. The hardstatus line can be changed with the ANSI Application Program Command (APC): ‘ESC_<string>ESC\’. As a convenience for xterm users the sequence ‘ESC]0..2;<string>^G’ is also accepted.

[Command] defhstatus [status]

(none) The hardstatus line that all new windows will get is set to status. This command is useful to make the hardstatus of every window display the window number or title or the like. status may contain the same directives as in the window messages, but the directive escape character is ‘^E’ (octal 005) instead of ‘%’. This was done to make a misinterpretation of program generated hardstatus lines impossible. If the parameter status is omitted, the current default string is displayed. Per default the hardstatus line of new windows is empty.

[Command] hstatus status

(none) Changes the current window’s hardstatus line to status.

10.9 Mousetrack

[Command] mousetrack [ on|off ]

(none) This command determines whether screen will watch for mouse clicks. When this command is enabled, regions that have been split in various ways can be selected by pointing to them with a mouse and left-clicking them. Without specifying on or off, the current state is displayed. The default state is determined by the defmousetrack command.

[Command] defmousetrack on|off

(none) This command determines the default state of the mousetrack command, currently defaulting of off.

11 Virtual Terminal

Each window in a screen session emulates a VT100 terminal, with some extra functions added. The VT100 emulator is hard-coded, no other terminal types can be emulated. The commands described here modify the terminal emulation.

11.1 Control Sequences

The following is a list of control sequences recognized by screen. ‘(V)’ and ‘(A)’ indicate VT100-specific and ANSI- or ISO-specific functions, respectively.

ESC E                           Next Line
ESC D                           Index
ESC M                           Reverse Index
ESC H                           Horizontal Tab Set
ESC Z                           Send VT100 Identification String
ESC 7                   (V)     Save Cursor and Attributes
ESC 8                   (V)     Restore Cursor and Attributes
ESC [s                  (A)     Save Cursor and Attributes
ESC [u                  (A)     Restore Cursor and Attributes
ESC c                           Reset to Initial State
ESC g                           Visual Bell
ESC Pn p                        Cursor Visibility (97801)
    Pn = 6                      Invisible
         7                      Visible
ESC =                   (V)     Application Keypad Mode
ESC >                   (V)     Numeric Keypad Mode
ESC # 8                 (V)     Fill Screen with E's
ESC \                   (A)     String Terminator
ESC ^(A)     Privacy Message String (Message Line)
ESC !                           Global Message String (Message Line)
ESC k                           Title Definition String
ESC P                   (A)     Device Control String
                                Outputs a string directly to the host
                                terminal without interpretation.
ESC _(A)     Application Program Command (Hardstatus)
ESC ] 0 ; string ^G     (A)     Operating System Command (Hardstatus, xterm
                                title hack)
ESC ] 83 ; cmd ^G       (A)     Execute screen command. This only works if
                                multi-user support is compiled into screen.
                                The pseudo-user ":window:" is used to check
                                the access control list. Use "addacl :window:
                                -rwx #?" to create a user with no rights and
                                allow only the needed commands.
Control-N               (A)     Lock Shift G1 (SO)
Control-O               (A)     Lock Shift G0 (SI)
ESC n                   (A)     Lock Shift G2
ESC o                   (A)     Lock Shift G3

ESC N                   (A)     Single Shift G2
ESC O                   (A)     Single Shift G3
ESC ( Pcs               (A)     Designate character set as G0
ESC ) Pcs               (A)     Designate character set as G1
ESC * Pcs               (A)     Designate character set as G2
ESC + Pcs               (A)     Designate character set as G3
ESC [ Pn ; Pn H                 Direct Cursor Addressing
ESC [ Pn ; Pn f                 same as above
ESC [ Pn J                      Erase in Display
Pn = None or 0            From Cursor to End of Screen
1                    From Beginning of Screen to Cursor
2                    Entire Screen
ESC [ Pn K                      Erase in Line
Pn = None or 0            From Cursor to End of Line
1                    From Beginning of Line to Cursor
2                    Entire Line
ESC [ Pn X                      Erase character
ESC [ Pn A                      Cursor Up
ESC [ Pn B                      Cursor Down
ESC [ Pn C                      Cursor Right
ESC [ Pn D                      Cursor Left
ESC [ Pn E                      Cursor next line
ESC [ Pn F                      Cursor previous line
ESC [ Pn G                      Cursor horizontal position
ESC [ Pn `                      same as above
ESC [ Pn d                      Cursor vertical position
ESC [ Ps ;...; Ps m             Select Graphic Rendition
Ps = None or 0            Default Rendition
1                    Bold
2            (A)     Faint
3            (A)     Standout Mode (ANSI: Italicized)
4                    Underlined
5                    Blinking
7                    Negative Image
22           (A)     Normal Intensity
23           (A)     Standout Mode off (ANSI: Italicized off)
24           (A)     Not Underlined
25           (A)     Not Blinking
27           (A)     Positive Image
30           (A)     Foreground Black
31           (A)     Foreground Red
32           (A)     Foreground Green
33           (A)     Foreground Yellow
34           (A)     Foreground Blue
35           (A)     Foreground Magenta
36           (A)     Foreground Cyan
37           (A)     Foreground White

           39           (A)     Foreground Default
           40           (A)     Background Black
           ...                  ...
           49           (A)     Background Default
ESC [ Pn g                      Tab Clear
      Pn = None or 0            Clear Tab at Current Position
           3                    Clear All Tabs
ESC [ Pn ; Pn r         (V)     Set Scrolling Region
ESC [ Pn I              (A)     Horizontal Tab
ESC [ Pn Z              (A)     Backward Tab
ESC [ Pn L              (A)     Insert Line
ESC [ Pn M              (A)     Delete Line
ESC [ Pn @              (A)     Insert Character
ESC [ Pn P              (A)     Delete Character
ESC [ Pn S                      Scroll Scrolling Region Up
ESC [ Pn T                      Scroll Scrolling Region Down
ESC [ Pn ^same as above
ESC [ Ps ;...; Ps h             Set Mode
ESC [ Ps ;...; Ps l             Reset Mode
      Ps = 4            (A)     Insert Mode
           20           (A)     `Automatic Linefeed' Mode.
           34                   Normal Cursor Visibility
           ?1           (V)     Application Cursor Keys
           ?3           (V)     Change Terminal Width to 132 columns
           ?5           (V)     Reverse Video
           ?6           (V)     `Origin' Mode
           ?7           (V)     `Wrap' Mode
           ?9                   X10 mouse tracking
           ?25          (V)     Visible Cursor
           ?47                  Alternate Screen (old xterm code)
           ?1000        (V)     VT200 mouse tracking
           ?1047                Alternate Screen (new xterm code)
           ?1049                Alternate Screen (new xterm code)
ESC [ 5 i               (A)     Start relay to printer (ANSI Media Copy)
ESC [ 4 i               (A)     Stop relay to printer (ANSI Media Copy)
ESC [ 8 ; Ph ; Pw t             Resize the window to `Ph' lines and
                                `Pw' columns (SunView special)
ESC [ c                         Send VT100 Identification String
ESC [ x                 (V)     Send Terminal Parameter Report
ESC [ > c                       Send Secondary Device Attributes String
ESC [ 6 n                       Send Cursor Position Report

11.2 Input Translation

In order to do a full VT100 emulation screen has to detect that a sequence of characters in the input stream was generated by a keypress on the user’s keyboard and insert the VT100 style escape sequence. Screen has a very flexible way of doing this by making it possible to map arbitrary commands on arbitrary sequences of characters. For standard VT100 emulation the command will always insert a string in the input buffer of the window (see also command stuff, see Section 12.2 [Paste], page 68). Because the sequences generated by a keypress can change after a reattach from a different terminal type, it is possible to bind commands to the termcap name of the keys. Screen will insert the correct binding after each reattach. See Section 14.5 [Bindkey], page 75, for further details on the syntax and examples.

Here is the table of the default key bindings. (A) means that the command is executed if the keyboard is switched into application mode.

Key name        Termcap name    Command
-----------------------------------------------------
Cursor up            ku         stuff \033[A
                                stuff \033OA      (A)
Cursor down          kd         stuff \033[B
                                stuff \033OB      (A)
Cursor right         kr         stuff \033[C
                                stuff \033OC      (A)
Cursor left          kl         stuff \033[D
                                stuff \033OD      (A)
Function key 0       k0         stuff \033[10~
Function key 1       k1         stuff \033OP
Function key 2       k2         stuff \033OQ
Function key 3       k3         stuff \033OR
Function key 4       k4         stuff \033OS
Function key 5       k5         stuff \033[15~
Function key 6       k6         stuff \033[17~
Function key 7       k7         stuff \033[18~
Function key 8       k8         stuff \033[19~
Function key 9       k9         stuff \033[20~
Function key 10      k;         stuff \033[21~
Function key 11      F1         stuff \033[23~
Function key 12      F2         stuff \033[24~
Home                 kh         stuff \033[1~
End                  kH         stuff \033[4~
Insert               kI         stuff \033[2~
Delete               kD         stuff \033[3~
Page up              kP         stuff \033[5~
Page down            kN         stuff \033[6~
Keypad 0             f0         stuff 0
                                stuff \033Op      (A)
Keypad 1             f1         stuff 1
                                stuff \033Oq      (A)
Keypad 2             f2         stuff 2
                                stuff \033Or      (A)

Keypad 3             f3         stuff 3
                                stuff \033Os      (A)
Keypad 4             f4         stuff 4
                                stuff \033Ot      (A)
Keypad 5             f5         stuff 5
                                stuff \033Ou      (A)
Keypad 6             f6         stuff 6
                                stuff \033Ov      (A)
Keypad 7             f7         stuff 7
                                stuff \033Ow      (A)
Keypad 8             f8         stuff 8
                                stuff \033Ox      (A)
Keypad 9             f9         stuff 9
                                stuff \033Oy      (A)
Keypad +             f+         stuff +
                                stuff \033Ok      (A)
Keypad -             f-         stuff -
                                stuff \033Om      (A)
Keypad *             f*         stuff *
                                stuff \033Oj      (A)
Keypad /             f/         stuff /
                                stuff \033Oo      (A)
Keypad =             fq         stuff =
                                stuff \033OX      (A)
Keypad .             f.         stuff .
                                stuff \033On      (A)
Keypad ,             f,         stuff ,
                                stuff \033Ol      (A)
Keypad enter         fe         stuff \015
                                stuff \033OM      (A)

11.3 Digraph

[Command] digraph [preset [unicode-value]]

(C-a C-v)

This command prompts the user for a digraph sequence. The next two characters typed are looked up in a builtin table and the resulting character is inserted in the input stream. For example, if the user enters ‘a"’, an a-umlaut will be inserted. If the first character entered is a 0 (zero), screen will treat the following characters (up to three) as an octal number instead. The optional argument preset is treated as user input, thus one can create an "umlaut" key. For example the command ‘bindkey ^K digraph '"'’ enables the user to generate an a-umlaut by typing ‘CTRL-K a’. When a non-zero unicode-value is specified, a new digraph is created with the specified preset. The digraph is unset if a zero value is provided for the unicode-value.

The following table is the builtin sequences.

          Sequence   Octal   Digraph   Unicode Equivalent

-----------------------------------------------
' ', ' '    160    (space)         U+00A0
'N', 'S'    160    (space)         U+00A0
'~', '!'    161       ¡            U+00A1
'!', '!'    161       ¡            U+00A1
'!', 'I'    161       ¡            U+00A1
'c', '|'    162  ¢  U+00A2
'c', 't'    162  ¢  U+00A2
'$', '$'    163  £  U+00A3
'P', 'd'    163  £  U+00A3
'o', 'x'    164                    U+00A4
'C', 'u'    164                    U+00A4
'C', 'u'    164                    U+00A4
'E', 'u'    164                    U+00A4
'Y', '-'    165  ¥  U+00A5
'Y', 'e'    165  ¥  U+00A5
'|', '|'    166       Š            U+00A6
'B', 'B'    166       Š            U+00A6
'p', 'a'    167  §  U+00A7
'S', 'E'    167  §  U+00A7
'"', '"'    168       š            U+00A8
''', ':'    168       š            U+00A8

'c', 'O'    169       \begin{matrix}c \\ ◯ \end{matrix}  U+00A9

'C', 'o'    169       \begin{matrix}c \\ ◯ \end{matrix}  U+00A9

'a', '-'    170       ^{\underline{a}}                U+00AA
'<', '<'    171       «            U+00AB
'-', ','    172  ¬  U+00AC
'N', 'O'    172  ¬  U+00AC
'-', '-'    173                   U+00AD

'r', 'O'    174       ◯^{R}         U+00AE

'R', 'g'    174       ◯^{R}         U+00AE

'-', '='    175       ¯ U+00AF
''', 'm'    175       ¯ U+00AF
'~', 'o'    176       ^{∘}  U+00B0
'D', 'G'    176       ^{∘}  U+00B0
'+', '-'    177  ±  U+00B1
'2', '2'    178       ^{2}  U+00B2
'2', 'S'    178       ^{2}  U+00B2
'3', '3'    179       ^{3}  U+00B3
'3', 'S'    179       ^{3}  U+00B3
''', '''    180       Ž            U+00B4
'j', 'u'    181  μ  U+00B5
'M', 'y'    181  μ  U+00B5
'p', 'p'    182  ¶  U+00B6
'P', 'I'    182  ¶  U+00B6
'~', '.'    183  ·  U+00B7
'.', 'M'    183  ·  U+00B7

',', ','    184       ž            U+00B8
''', ','    184       ž            U+00B8
'1', '1'    185       ^{1}  U+00B9
'1', 'S'    185       ^{1}  U+00B9
'o', '-'    186       ^{\underline{o}}                U+00BA
'>', '>'    187       »            U+00BB
'1', '4'    188       Œ            U+00BC
'1', '2'    189       œ            U+00BD
'3', '4'    190       Ÿ            U+00BE
'~', '?'    191       ¿            U+00BF
'?', '?'    191       ¿            U+00BF
'?', 'I'    191       ¿            U+00BF
'A', '`'    192       À            U+00C0
'A', '!'    192       À            U+00C0
'A', '''    193       Á            U+00C1
'A', '^'    194       ^A            U+00C2
'A', '>'    194       ^A            U+00C2
'A', '~'    195       ~ A            U+00C3
'A', '?'    195       ~ A            U+00C3
'A', '"'    196       Ä            U+00C4
'A', ':'    196       Ä            U+00C4
'A', '@'    197       Å            U+00C5
'A', 'A'    197       Å            U+00C5
'A', 'E'    198       Æ            U+00C6
'C', ','    199       Ç            U+00C7
'E', '`'    200       È            U+00C8
'E', '!'    200       È            U+00C8
'E', '''    201       É            U+00C9
'E', '^'    202       ^E            U+00CA
'E', '>'    202       ^E            U+00CA
'E', '"'    203       Ë            U+00CB
'E', ':'    203       Ë            U+00CB
'I', '`'    204       Ì            U+00CC
'I', '!'    204       Ì            U+00CC
'I', '''    205       Í            U+00CD
'I', '^'    206       ^I            U+00CE
'I', '>'    206       ^I            U+00CE
'I', '"'    207       Ï            U+00CF
'I', ':'    207       Ï            U+00CF
'D', '-'    208       Ð            U+00D0
'N', '~'    209       ~ N            U+00D1
'N', '?'    209       ~ N            U+00D1
'O', '`'    210       Ò            U+00D2
'O', '!'    210       Ò            U+00D2
'O', '''    211       Ó            U+00D3
'O', '^'    212       ^O            U+00D4
'O', '>'    212       ^O            U+00D4

'O', '~'    213       ~ O            U+00D5
'O', '?'    213       ~ O            U+00D5
'O', '"'    214       Ö            U+00D6
'O', ':'    214       Ö            U+00D6
'/', '\'    215  ×  U+00D7
'*', 'x'    215  ×  U+00D7
'O', '/'    216       Ø            U+00D8
'U', '`'    217       Ù            U+00D9
'U', '!'    217       Ù            U+00D9
'U', '''    218       Ú            U+00DA
'U', '^'    219       ^U            U+00DB
'U', '>'    219       ^U            U+00DB
'U', '"'    220       Ü            U+00DC
'U', ':'    220       Ü            U+00DC
'Y', '''    221       Ý            U+00DD
'I', 'p'    222       Þ            U+00DE
'T', 'H'    222       Þ            U+00DE
's', 's'    223       ß            U+00DF
's', '"'    223       ß            U+00DF
'a', '`'    224       à            U+00E0
'a', '!'    224       à            U+00E0
'a', '''    225       á            U+00E1
'a', '^'    226       ^a            U+00E2
'a', '>'    226       ^a            U+00E2
'a', '~'    227       ~ a            U+00E3
'a', '?'    227       ~ a            U+00E3
'a', '"'    228       ä            U+00E4
'a', ':'    228       ä            U+00E4
'a', 'a'    229       å            U+00E5
'a', 'e'    230       æ            U+00E6
'c', ','    231       ç            U+00E7
'e', '`'    232       è            U+00E8
'e', '!'    232       è            U+00E8
'e', '''    233       é            U+00E9
'e', '^'    234       ^e            U+00EA
'e', '>'    234       ^e            U+00EA
'e', '"'    235       ë            U+00EB
'e', ':'    235       ë            U+00EB
'i', '`'    236       ì            U+00EC
'i', '!'    236       ì            U+00EC
'i', '''    237       í            U+00ED
'i', '^'    238       ^ı            U+00EE
'i', '>'    238       ^ı            U+00EE
'i', '"'    239       ï            U+00EF
'i', ':'    239       ï            U+00EF
'd', '-'    240       ð            U+00F0
'n', '~'    241       ~ n            U+00F1

'n', '?'    241       ~ n            U+00F1
'o', '`'    242       ò            U+00F2
'o', '!'    242       ò            U+00F2
'o', '''    243       ó            U+00F3
'o', '^'    244       ^o            U+00F4
'o', '>'    244       ^o            U+00F4
'o', '~'    245       ~ o            U+00F5
'o', '?'    245       ~ o            U+00F5
'o', '"'    246       ö            U+00F6
'o', ':'    246       ö            U+00F6
':', '-'    247  ÷  U+00F7
'o', '/'    248       ø            U+00F8
'u', '`'    249       ù            U+00F9
'u', '!'    249       ù            U+00F9
'u', '''    250       ú            U+00FA
'u', '^'    251       ^u            U+00FB
'u', '>'    251       ^u            U+00FB
'u', '"'    252       ü            U+00FC
'u', ':'    252       ü            U+00FC
'y', '''    253       ý            U+00FD
'i', 'p'    254       þ            U+00FE
't', 'h'    254       þ            U+00FE
'y', '"'    255       ÿ            U+00FF
'y', ':'    255       ÿ            U+00FF
'"', '['    196       Ä            U+00C4
'"', '\'    214       Ö            U+00D6
'"', ']'    220       Ü            U+00DC
'"', '{'    228       ä            U+00E4
'"', '|'    246       ö            U+00F6
'"', '}'    252       ü            U+00FC
'"', '~'    223       ß            U+00DF

11.4 Bell

[Command] bell_msg [message]

(none) When a bell character is sent to a background window, screen displays a notification in the message line. The notification message can be re-defined by this command. Each occurrence of ‘%’ in message is replaced by the number of the window to which a bell has been sent, and each occurrence of ‘^G’ is replaced by the definition for bell in your termcap (usually an audible bell). The default message is

'Bell in window %n'

An empty message can be supplied to the bell_msg command to suppress output of a message line (bell_msg ""). Without a parameter, the current message is shown.

[Command] vbell [state]

(C-a C-g)

Sets or toggles the visual bell setting for the current window. If vbell is switched to ‘on’, but your terminal does not support a visual bell, the visual bell message is displayed in the status line when the bell character is received. Visual bell support of a terminal is defined by the termcap variable vb. See Section 11.4 [Bell], page 59, for more information on visual bells. The equivalent terminfo capability is flash. Per default, vbell is ‘off’, thus the audible bell is used.

[Command] vbell_msg [message]

(none) Sets the visual bell message. Message is printed to the status line if the window receives a bell character (^G), vbell is set to ‘on’ and the terminal does not support a visual bell. The default message is ‘Wuff, Wuff!!’. Without a parameter, the current message is shown.

[Command] vbellwait sec

(none) Define a delay in seconds after each display of screen ’s visual bell message. The default is 1 second.

11.5 Clear

[Command] clear

(C-a C)

Clears the screen and saves its contents to the scrollback buffer.

11.6 Info

[Command] info

(C-a i, C-a C-i)

Uses the message line to display some information about the current window: the cursor position in the form ‘(column,row)’ starting with ‘(1,1)’, the terminal width and height plus the size of the scrollback buffer in lines, like in ‘(80,24)+50’, the current state of window XON/XOFF flow control is shown like this (see Chapter 15 [Flow Control], page 77):

       +flow     automatic flow control, currently on.
       -flow     automatic flow control, currently off.
       +(+)flow  flow control enabled. Agrees with automatic control.
       -(+)flow  flow control disabled. Disagrees with automatic control.
       +(-)flow  flow control enabled. Disagrees with automatic control.
       -(-)flow  flow control disabled. Agrees with automatic control.

The current line wrap setting (‘+wrap’ indicates enabled, ‘-wrap’ not) is also shown. The flags ‘ins’, ‘org’, ‘app’, ‘log’, ‘mon’ and ‘nored’ are displayed when the window is in insert mode, origin mode, application-keypad mode, has output logging, activity monitoring or partial redraw enabled.

The currently active character set (‘G0’, ‘G1’, ‘G2’, or ‘G3’), and in square brackets the terminal character sets that are currently designated as ‘G0’ through ‘G3’. If the window is in UTF-8 mode, the string ‘UTF-8’ is shown instead. Additional modes depending on the type of the window are displayed at the end of the status line (see Section 6.6 [Window Types], page 29). If the state machine of the terminal emulator is in a non-default state, the info line is started with a string identifying the current state. For system information use time.

[Command] dinfo

(none) Show what screen thinks about your terminal. Useful if you want to know why features like color or the alternate charset don’t work.

11.7 Redisplay

[Command] allpartial state

(none) If set to on, only the current cursor line is refreshed on window change. This affects all windows and is useful for slow terminal lines. The previous setting of full/partial refresh for each window is restored with allpartial off. This is a global flag that immediately takes effect on all windows overriding the partial settings. It does not change the default redraw behavior of newly created windows.

[Command] altscreen state

(none) If set to on, "alternate screen" support is enabled in virtual terminals, just like in xterm. Initial setting is ‘off’.

[Command] partial state

(none) Defines whether the display should be refreshed (as with redisplay) after switching to the current window. This command only affects the current window. To immediately affect all windows use the allpartial command. Default is ‘off’, of course. This default is fixed, as there is currently no defpartial command.

[Command] redisplay
(C-a l, C-a C-l)

Redisplay the current window. Needed to get a full redisplay in partial redraw mode.

11.8 Wrap

[Command] wrap [ on | off ]

(C-a r, C-a C-r)

Sets the line-wrap setting for the current window. When line-wrap is on, the second consecutive printable character output at the last column of a line will wrap to the start of the following line. As an added feature, backspace (^H) will also wrap through the left margin to the previous line. Default is ‘on’. Without any options, the state of wrap is toggled.

[Command] defwrap state

(none) Same as the wrap command except that the default setting for new windows is changed. Initially line-wrap is on and can be toggled with the wrap command (C-a r) or by means of "C-a : wrap on|off".

11.9 Reset

[Command] reset

(C-a Z)

Reset the virtual terminal to its “power-on” values. Useful when strange settings (like scroll regions or graphics character set) are left over from an application.

11.10 Window Size

[Command] width [-w|-d] [cols [lines]]

(C-a W)

Toggle the window width between 80 and 132 columns, or set it to cols columns if an argument is specified. This requires a capable terminal and the termcap entries ‘Z0’ and ‘Z1’. See the termcap command (see Chapter 16 [Termcap], page 79), for more information. You can also specify a height if you want to change both values. The -w option tells screen to leave the display size unchanged and just set the window size, -d vice versa.

[Command] height [-w|-d] [lines [cols]]

(none) Set the display height to a specified number of lines. When no argument is given it toggles between 24 and 42 lines display.

11.11 Character Processing

[Command] c1 [state]

(none) Change c1 code processing. ‘c1 on’ tells screen to treat the input characters between 128 and 159 as control functions. Such an 8-bit code is normally the same as ESC followed by the corresponding 7-bit code. The default setting is to process c1 codes and can be changed with the ‘defc1’ command. Users with fonts that have usable characters in the c1 positions may want to turn this off.

[Command] gr [state]

(none) Turn GR charset switching on/off. Whenever screen sees an input char with an 8th bit set, it will use the charset stored in the GR slot and print the character with the 8th bit stripped. The default (see also ‘defgr’) is not to process GR switching because otherwise the ISO88591 charset would not work.

[Command] bce [state]

(none) Change background-color-erase setting. If ‘bce’ is set to on, all characters cleared by an erase/insert/scroll/clear operation will be displayed in the current background color. Otherwise the default background color is used.

[Command] encoding enc [denc]

(none) Tell screen how to interpret the input/output. The first argument sets the encoding of the current window. Each window can emulate a different encoding. The optional second parameter overwrites the encoding of the connected terminal. It should never be needed as screen uses the locale setting to detect the encoding. There is also a way to select a terminal encoding depending on the terminal type by using the ‘KJ’ termcap entry. See Section 16.5 [Special Capabilities], page 82.

Supported encodings are eucJP, SJIS, eucKR, eucCN, Big5, GBK, KOI8-R, CP1251,
UTF-8, ISO8859-2, ISO8859-3, ISO8859-4, ISO8859-5, ISO8859-6, ISO8859-7,
ISO8859-8, ISO8859-9, ISO8859-10, ISO8859-15, jis.

See also ‘defencoding’, which changes the default setting of a new window.

[Command] charset set

(none) Change the current character set slot designation and charset mapping. The first four character of set are treated as charset designators while the fifth and sixth character must be in range ‘0’ to ‘3’ and set the GL/GR charset mapping. On every position a ‘.’ may be used to indicate that the corresponding charset/mapping should not be changed (set is padded to six characters internally by appending ‘.’ chars). New windows have ‘BBBB02’ as default charset, unless a ‘encoding’ command is active. The current setting can be viewed with the Section 11.6 [Info], page 60, command.

[Command] utf8 [state [dstate]]

(none) Change the encoding used in the current window. If utf8 is enabled, the strings sent to the window will be UTF-8 encoded and vice versa. Omitting the parameter toggles the setting. If a second parameter is given, the display’s encoding is also changed (this should rather be done with screen’s ‘-U’ option). See also ‘defutf8’, which changes the default setting of a new window.

[Command] defc1 state

(none) Same as the ‘c1’ command except that the default setting for new windows is changed. Initial setting is ‘on’.

[Command] defgr state

(none) Same as the ‘gr’ command except that the default setting for new windows is changed. Initial setting is ‘off’.

[Command] defbce state

(none) Same as the ‘bce’ command except that the default setting for new windows is changed. Initial setting is ‘off’.

[Command] defencoding enc

(none) Same as the ‘encoding’ command except that the default setting for new windows is changed. Initial setting is the encoding taken from the terminal.

[Command] defcharset [set]

(none) Like the ‘charset’ command except that the default setting for new windows is changed. Shows current default if called without argument.

[Command] defutf8 state

(none) Same as the ‘utf8’ command except that the default setting for new windows is changed. Initial setting is on if screen was started with ‘-U’, otherwise off.

[Command] cjkwidth [state]

(none) Toggle how ambiguoous characters are treated. If cjkwidth is on screen interprets them as double (full) width characters. If off then they are seen as one cell (half) width characters.

12 Copy and Paste

For those confined to a hardware terminal, these commands provide a cut and paste facility more powerful than those provided by most windowing systems.

12.1 Copying

[Command] copy

(C-a [, C-a C-[, C-a ESC)

Enter copy/scrollback mode. This allows you to copy text from the current window and its history into the paste buffer. In this mode a vi-like full screen editor is active, with controls as outlined below.

12.1.1 CR/LF

[Command] crlf [state]

(none) This affects the copying of text regions with the copy command. If it is set to ‘on’, lines will be separated by the two character sequence ‘CR’/‘LF’. Otherwise only ‘LF’ is used. crlf is off by default. When no parameter is given, the state is toggled.

12.1.2 Scrollback

To access and use the contents in the scrollback buffer, use the copy command. See Section 12.1 [Copy], page 65.

[Command] defscrollback num

(none) Same as the scrollback command except that the default setting for new windows is changed. Defaults to 100.

[Command] scrollback num

(none) Set the size of the scrollback buffer for the current window to num lines. The default scrollback is 100 lines. Use info to view the current setting.

[Command] compacthist [state]

(none) This tells screen whether to suppress trailing blank lines when scrolling up text into the history buffer. Turn compacting ‘on’ to hold more useful lines in your scrollback buffer.

12.1.3 Markkeys

[Command] markkeys string

(none) This is a method of changing the keymap used for copy/history mode. The string is made up of oldchar=newchar pairs which are separated by ‘:’. Example: The command markkeys h=^B:l=^F:$=^E would set some keys to be more familiar to emacs users. If your terminal sends characters, that cause you to abort copy mode, then this command may help by binding these characters to do nothing. The no-op character is ‘@’ and is used like this: markkeys @=L=H if you do not want to use the ‘H’ or ‘L’ commands any longer. As shown in this example, multiple keys can be assigned to one function in a single statement.

12.1.4 Movement Keys

h, C-h, or left arrow move the cursor left. j, C-n, or down arrow move the cursor down. k, C-p, or up arrow move the cursor up. l (’el’), or right arrow move the cursor right. 0 (zero) or C-a move to the leftmost column. + and - move the cursor to the leftmost column of the next or previous line. H, M and L move the cursor to the leftmost column of the top, center or bottom line of the window. | moves to the specified absolute column. g or home moves to the beginning of the buffer. G or end moves to the specified absolute line (default: end of buffer). % jumps to the specified percentage of the buffer. ^or $ move to the first or last non-whitespace character on the line. w, b, and e move the cursor word by word. B, E move the cursor WORD by WORD (as in vi). f/F, t/T move the cursor forward/backward to the next occurence of the target. (eg, ’3fy’ will move the cursor to the 3rd ’y’ to the right.) ; and , Repeat the last f/F/t/T command in the same/opposite direction. C-e and C-y scroll the display up/down by one line while preserving the cursor position. C-u and C-d scroll the display up/down by the specified amount of lines while preserving the cursor position. (Default: half screenful). C-b and C-f move the cursor up/down a full screen.

Note that Emacs-style movement keys can be specified by a .screenrc command. (markkeys "h=^B:l=^F:$=^E") There is no simple method for a full emacs-style keymap, however, as this involves multi-character codes.

12.1.5 Marking

The copy range is specified by setting two marks. The text between these marks will be highlighted. Press: space or enter to set the first or second mark respectively. If mousetrack is set to on, marks can also be set using left mouse click. Y and y can be used to mark one whole line or to mark from start of line. W marks exactly one word.

12.1.6 Repeat Count

Any command in copy mode can be prefixed with a number (by pressing digits 0...9) which is taken as a repeat count. Example:

     C-a C-[ H 10 j 5 Y

will copy lines 11 to 15 into the paste buffer.

12.1.7 Searching

/ vi-like search forward. ? vi-like search backward. C-a s emacs style incremental search forward. C-r emacs style reverse i-search.

[Command] ignorecase [on|off]

(none)

Tell screen to ignore the case of characters in searches. Default is off. Without any options, the state of ignorecase is toggled.

n Repeat search in forward direction. N Repeat search in backward direction.

12.1.8 Specials

There are, however, some keys that act differently here from in vi. Vi does not allow to yank rectangular blocks of text, but screen does. Press: c or C to set the left or right margin respectively. If no repeat count is given, both default to the current cursor position. Example: Try this on a rather full text screen:

     C-a [ M 20 l SPACE c 10 l 5 j C SPACE.

This moves one to the middle line of the screen, moves in 20 columns left, marks the beginning of the paste buffer, sets the left column, moves 5 columns down, sets the right column, and then marks the end of the paste buffer. Now try:

     C-a [ M 20 l SPACE 10 l 5 j SPACE

and notice the difference in the amount of text copied. J joins lines. It toggles between 4 modes: lines separated by a newline character (012), lines glued seamless, lines separated by a single space or comma separated lines. Note that you can prepend the newline character with a carriage return character, by issuing a set crlf

on.

v or V is for all the vi users who use :set numbers - it toggles the left margin between column 9 and 1. a before the final space key turns on append mode. Thus the contents of the paste buffer will not be overwritten, but appended to. A turns on append mode and sets a (second) mark. > sets the (second) mark and writes the contents of the paste buffer to the screen-exchange file (/tmp/screen-exchange per default) once copy-mode is finished. See Section 12.4 [Screen Exchange], page 69. This example demonstrates how to dump the whole scrollback buffer to that file:

C-a [ g SPACE G $ >.

C-g gives information about the current line and column. x or o (’oh’) exchanges the first mark and the current cursor position. You can use this to adjust an already placed mark. C-l (’el’) will redraw the screen. @ does nothing. Absolutely nothing. Does not even exit copy mode. All keys not described here exit copy mode.

12.2 Paste

[Command] paste [registers [destination]]

(C-a ], C-a C-])

Write the (concatenated) contents of the specified registers to the stdin stream of the current window. The register ‘.’ is treated as the paste buffer. If no parameter is specified the user is prompted to enter a single register. The paste buffer can be filled with the copy, history and readbuf commands. Other registers can be filled with the register, readreg and paste commands. If paste is called with a second argument, the contents of the specified registers is pasted into the named destination register rather than the window. If ‘.’ is used as the second argument, the display’s paste buffer is the destination. Note, that paste uses a wide variety of resources: Usually both, a current window and a current display are required. But whenever a second argument is specified no current window is needed. When the source specification only contains registers (not the paste buffer) then there need not be a current display (terminal attached), as the registers are a global resource. The paste buffer exists once for every user.

[Command] stuff [string]

(none) Stuff the string string in the input buffer of the current window. This is like the paste command, but with much less overhead. Without a paramter, screen will prompt for a string to stuff. You cannot paste large buffers with the stuff command. It is most useful for key bindings. See Section 14.5 [Bindkey], page 75.

[Command] pastefont [state]

Tell screen to include font information in the paste buffer. The default is not to do so. This command is especially useful for multi character fonts like kanji.

[Command] slowpaste msec

[Command] defslowpaste msec

(none) Define the speed text is inserted in the current window by the paste command. If the slowpaste value is nonzero text is written character by character. screen will pause for msec milliseconds after each write to allow the application to process the input. only use slowpaste if your underlying system exposes flow control problems while pasting large amounts of text. defslowpaste specifies the default for new windows.

[Command] readreg [-e encoding] [register [filename]]

(none) Does one of two things, dependent on number of arguments: with zero or one arguments it it duplicates the paste buffer contents into the register specified or entered at the prompt. With two arguments it reads the contents of the named file into the register, just as readbuf reads the screen-exchange file into the paste buffer. You can tell screen the encoding of the file via the -e option. The following example will paste the system’s password file into the screen window (using register p, where a copy remains):

     C-a : readreg p /etc/passwd
     C-a : paste p

12.3 Registers

[Command] copy_reg [key]

(none) Removed. Use readreg instead.

[Command] ins_reg [key]

(none) Removed. Use paste instead.

[Command] process [key]

(none) Stuff the contents of the specified register into the screen input queue. If no argument is given you are prompted for a register name. The text is parsed as if it had been typed in from the user’s keyboard. This command can be used to bind multiple actions to a single key.

[Command] register [-e encoding] key string

(none) Save the specified string to the register key. The encoding of the string can be specified via the -e option.

12.4 Screen Exchange

[Command] bufferfile [exchange-file]

(none) Change the filename used for reading and writing with the paste buffer. If the exchange-file parameter is omitted, screen reverts to the default of /tmp/screen-exchange. The following example will paste the system’s password file into the screen window (using the paste buffer, where a copy remains):

     C-a : bufferfile /etc/passwd
     C-a < C-a ]
     C-a : bufferfile

[Command] readbuf [-e encoding] [filename]
(C-a <)

Reads the contents of the specified file into the paste buffer. You can tell screen the encoding of the file via the -e option. If no file is specified, the screen-exchange filename is used.

[Command] removebuf
(C-a =)

Unlinks the screen-exchange file.

[Command] writebuf [-e encoding] [filename]
(C-a >)

Writes the contents of the paste buffer to the specified file, or the public accessible screen-exchange file if no filename is given. This is thought of as a primitive means of communication between screen users on the same host. If an encoding is specified the paste buffer is recoded on the fly to match the encoding. See also C-a ESC (see Section 12.1 [Copy], page 65).

12.5 History

[Command] history

(C-a {, C-a })

Usually users work with a shell that allows easy access to previous commands. For example, csh has the command !! to repeat the last command executed. screen provides a primitive way of recalling “the command that started …”: You just type the first letter of that command, then hit C-a { and screen tries to find a previous line that matches with the prompt character to the left of the cursor. This line is pasted into this window’s input queue. Thus you have a crude command history (made up by the visible window and its scrollback buffer).

13 Subprocess Execution

Control Input or Output of a window by another filter process. Use with care!

13.1 Exec

[Command] exec [[fdpat] newcommand [args ... ]]

(none) Run a unix subprocess (specified by an executable path newcommand and its optional arguments) in the current window. The flow of data between newcommands stdin/stdout/stderr, the process originally started (let us call it "application-process") and screen itself (window) is controlled by the file descriptor pattern fdpat. This pattern is basically a three character sequence representing stdin, stdout and stderr of newcommand. A dot (.) connects the file descriptor to screen. An exclamation mark (!) causes the file descriptor to be connected to the application-process. A colon (:) combines both. User input will go to newcommand unless newcommand receives the applicationprocess’ output (fdpats first character is ‘!’ or ‘:’) or a pipe symbol (‘|’) is added to the end of fdpat. Invoking exec without arguments shows name and arguments of the currently running subprocess in this window. Only one subprocess can be running per window. When a subprocess is running the kill command will affect it instead of the windows process. Only one subprocess a time can be running in each window. Refer to the postscript file doc/fdpat.ps for a confusing illustration of all 21 possible combinations. Each drawing shows the digits 2, 1, 0 representing the three file descriptors of newcommand. The box marked ‘W’ is usual pty that has the application-process on its slave side. The box marked ‘P’ is the secondary pty that now has screen at its master side.

13.2 Using Exec

Abbreviations:

• Whitespace between the word ‘exec’ and fdpat and the command name can be omitted.

• Trailing dots and a fdpat consisting only of dots can be omitted.

• A simple ‘|’ is synonymous for the ‘!..|’ pattern.

• The word ‘exec’ can be omitted when the ‘|’ abbreviation is used.

• The word ‘exec’ can always be replaced by leading ‘!’.

Examples:

!/bin/sh
exec /bin/sh
exec ... /bin/sh

All of the above are equivalent. Creates another shell in the same window, while the original shell is still running. Output of both shells is displayed and user input is sent to the new /bin/sh.

!!stty 19200
exec!stty 19200
exec !.. stty 19200

All of the above are equivalent. Set the speed of the window’s tty. If your stty command operates on stdout, then add another ‘!’. This is a useful command, when a screen window is directly connected to a serial line that needs to be configured.

|less
exec !..| less

Both are equivalent. This adds a pager to the window output. The special character ‘|’ is needed to give the user control over the pager although it gets its input from the window’s process. This works, because ‘less’ listens on stderr (a behavior that screen would not expect without the ‘|’) when its stdin is not a tty. Less versions newer than 177 fail miserably here; good old pg still works.

!:sed -n s/.*Error.*/\007/p

Sends window output to both, the user and the sed command. The sed inserts an additional bell character (oct. 007) to the window output seen by screen. This will cause ’Bell in window x’ messages, whenever the string ‘Error’ appears in the window.

14 Key Binding

You may disagree with some of the default bindings (I know I do). The bind command allows you to redefine them to suit your preferences.

14.1 The bind command

[Command] bind [-c class] key [command [args]]

(none) Bind a command to a key. The key argument is either a single character, a twocharacter sequence of the form ‘^x’ (meaning C-x), a backslash followed by an octal number (specifying the ASCII code of the character), or a backslash followed by a second character, such as ‘\^’ or ‘\\’. The argument can also be quoted, if you like. If no further argument is given, any previously established binding for this key is removed. The command argument can be any command (see [Command Index], page 115).

If a command class is specified via the -c option, the key is bound for the specified class. Use the command command to activate a class. Command classes can be used to create multiple command keys or multi-character bindings.

By default, most suitable commands are bound to one or more keys (see Section 5.1 [Default Key Bindings], page 11); for instance, the command to create a new window is bound to C-c and c. The bind command can be used to redefine the key bindings and to define new bindings.

[Command] unbindall

(none) Unbind all the bindings. This can be useful when screen is used solely for its detaching abilities, such as when letting a console application run as a daemon. If, for some reason, it is necessary to bind commands after this, use ’screen -X’.

14.2 Examples of the bind command

Some examples:

bind ' ' windows
bind ^f screen telnet foobar
bind \033 screen -ln -t root -h 1000 9 su

would bind the space key to the command that displays a list of windows (so that the command usually invoked by C-a C-w would also be available as C-a space), bind C-f to the command “create a window with a TELNET connection to foobar”, and bind ESC to the command that creates an non-login window with title ‘root’ in slot #9, with a superuser shell and a scrollback buffer of 1000 lines.

bind -c demo1 0 select 10
bind -c demo1 1 select 11
bind -c demo1 2 select 12
bindkey "^B" command -c demo1

makes C-b 0 select window 10, C-b 1 window 11, etc.

bind -c demo2 0 select 10
bind -c demo2 1 select 11
bind -c demo2 2 select 12
bind - command -c demo2

makes C-a - 0 select window 10, C-a - 1 window 11, etc.

14.3 Command Character

[Command] escape xy

(none) Set the command character to x and the character generating a literal command character (by triggering the meta command) to y (similar to the ‘-e’ option). Each argument is either a single character, a two-character sequence of the form ‘^x’ (meaning C-x), a backslash followed by an octal number (specifying the ASCII code of the character), or a backslash followed by a second character, such as ‘\^’ or ‘\\’. The default is ‘^Aa’, but ‘``’ is recommended by one of the authors.

[Command] defescape xy

(none) Set the default command characters. This is equivalent to the command escape except that it is useful for multiuser sessions only. In a multiuser session escape changes the command character of the calling user, where defescape changes the default command characters for users that will be added later.

[Command] meta

(C-a a)

Send the command character (C-a) to the process in the current window. The keystroke for this command is the second parameter to the ‘-e’ command line switch (see Chapter 3 [Invoking Screen], page 5), or the escape .screenrc directive.

[Command] command [-c class]

(none) This command has the same effect as typing the screen escape character (C-a). It is probably only useful for key bindings. If the ‘-c’ option is given, select the specified command class. See Section 14.1 [Bind], page 73, See Section 14.5 [Bindkey], page 75.

14.4 Help

[Command] help

(C-a ?)

Displays a help screen showing you all the key bindings. The first pages list all the internal commands followed by their bindings. Subsequent pages will display the custom commands, one command per key. Press space when you’re done reading each page, or return to exit early. All other characters are ignored. If the ‘-c’ option is given, display all bound commands for the specified command class. See Section 5.1 [Default Key Bindings], page 11.

14.5 Bindkey

[Command] bindkey [opts] [string [cmd args]]

(none) This command manages screen’s input translation tables. Every entry in one of the tables tells screen how to react if a certain sequence of characters is encountered. There are three tables: one that should contain actions programmed by the user, one for the default actions used for terminal emulation and one for screen’s copy mode to do cursor movement. See Section 11.2 [Input Translation], page 53, for a list of default key bindings. If the ‘-d’ option is given, bindkey modifies the default table, ‘-m’ changes the copy mode table and with neither option the user table is selected. The argument ‘string’ is the sequence of characters to which an action is bound. This can either be a fixed string or a termcap keyboard capability name (selectable with the ‘-k’ option). Some keys on a VT100 terminal can send a different string if application mode is turned on (e.g. the cursor keys). Such keys have two entries in the translation table. You can select the application mode entry by specifying the ‘-a’ option. The ‘-t’ option tells screen not to do inter-character timing. One cannot turn off the timing if a termcap capability is used. ‘cmd’ can be any of screen’s commands with an arbitrary number of ‘args’. If ‘cmd’ is omitted the key-binding is removed from the table.

14.6 Bindkey Examples

Here are some examples of keyboard bindings:

bindkey -d

Show all of the default key bindings. The application mode entries are marked with [A].

bindkey -k k1 select 1

Make the "F1" key switch to window one.

bindkey -t foo stuff barfoo

Make ‘foo’ an abbreviation of the word ‘barfoo’. Timeout is disabled so that users can type slowly.

bindkey "\024" mapdefault

This key-binding makes ‘C-t’ an escape character for key-bindings. If you did the above ‘stuff barfoo’ binding, you can enter the word ‘foo’ by typing ‘C-t foo’. If you want to insert a ‘C-t’ you have to press the key twice (i.e., escape the escape binding).

bindkey -k F1 command

Make the F11 (not F1!) key an alternative screen escape (besides ‘C-a’).

14.7 Bindkey Control

[Command] mapdefault

(none) Tell screen that the next input character should only be looked up in the default bindkey table.

[Command] mapnotnext

(none) Like mapdefault, but don’t even look in the default bindkey table.

[Command] maptimeout n

(none) Set the inter-character timer for input sequence detection to a timeout of n ms. The default timeout is 300ms. Maptimeout with no arguments shows the current setting.

15 Flow Control

screen can trap flow control characters or pass them to the program, as you see fit. This is useful when your terminal wants to use XON/XOFF flow control and you are running a program which wants to use ^S/^Q for other purposes (i.e. emacs).

15.1 About screen flow control settings

Each window has a flow-control setting that determines how screen deals with the XON and XOFF characters (and perhaps the interrupt character). When flow-control is turned off, screen ignores the XON and XOFF characters, which allows the user to send them to the current program by simply typing them (useful for the emacs editor, for instance). The trade-off is that it will take longer for output from a “normal” program to pause in response to an XOFF. With flow-control turned on, XON and XOFF characters are used to immediately pause the output of the current window. You can still send these characters to the current program, but you must use the appropriate two-character screen commands (typically C-a q (xon) and C-a s (xoff)). The xon/xoff commands are also useful for typing C-s and C-q past a terminal that intercepts these characters.

Each window has an initial flow-control value set with either the ‘-f’ option or the defflow command. By default the windows are set to automatic flow-switching. It can then be toggled between the three states ’fixed on’, ’fixed off’ and ’automatic’ interactively with the flow command bound to C-a f.

The automatic flow-switching mode deals with flow control using the TIOCPKT mode (like rlogin does). If the tty driver does not support TIOCPKT, screen tries to determine the right mode based on the current setting of the application keypad — when it is enabled, flow-control is turned off and visa versa. Of course, you can still manipulate flow-control manually when needed.

If you’re running with flow-control enabled and find that pressing the interrupt key (usually C-c) does not interrupt the display until another 6-8 lines have scrolled by, try running screen with the ‘interrupt’ option (add the ‘interrupt’ flag to the flow command in your .screenrc, or use the ‘-i’ command-line option). This causes the output that screen has accumulated from the interrupted program to be flushed. One disadvantage is that the virtual terminal’s memory contains the non-flushed version of the output, which in rare cases can cause minor inaccuracies in the output. For example, if you switch screens and return, or update the screen with C-a l you would see the version of the output you would have gotten without ‘interrupt’ being on. Also, you might need to turn off flow-control (or use auto-flow mode to turn it off automatically) when running a program that expects you to type the interrupt character as input, as the ‘interrupt’ parameter only takes effect when flow-control is enabled. If your program’s output is interrupted by mistake, a simple refresh of the screen with C-a l will restore it. Give each mode a try, and use whichever mode you find more comfortable.

15.2 Flow

[Command] defflow fstate [interrupt]

(none)

Same as the flow command except that the default setting for new windows is changed. Initial setting is ‘auto’. Specifying flow auto interrupt has the same effect as the command-line options ‘-fa’ and ‘-i’. Note that if ‘interrupt’ is enabled, all existing displays are changed immediately to forward interrupt signals.

[Command] flow [fstate]

(C-a f, C-a C-f)

Sets the flow-control mode for this window to fstate, which can be ‘on’, ‘off’ or ‘auto’. Without parameters it cycles the current window’s flow-control setting. Default is set by ‘defflow’.

15.3 XON and XOFF

[Command] xon

(C-a q, C-a C-q)

Send a ^Q (ASCII XON) to the program in the current window. Redundant if flow control is set to ‘off’ or ‘auto’.

[Command] xoff

(C-a s, C-a C-s)

Send a ^S (ASCII XOFF) to the program in the current window.

16 Termcap

Screen demands the most out of your terminal so that it can perform its VT100 emulation most efficiently. These functions provide means for tweaking the termcap entries for both your physical terminal and the one simulated by screen.

16.1 Choosing the termcap entry for a window

Usually screen tries to emulate as much of the VT100/ANSI standard as possible. But if your terminal lacks certain capabilities the emulation may not be complete. In these cases screen has to tell the applications that some of the features are missing. This is no problem on machines using termcap, because screen can use the $TERMCAP variable to customize the standard screen termcap.

But if you do a rlogin on another machine or your machine supports only terminfo this method fails. Because of this screen offers a way to deal with these cases. Here is how it works:

When screen tries to figure out a terminal name for itself, it first looks for an entry named screen.term, where term is the contents of your $TERM variable. If no such entry exists, screen tries ‘screen’ (or ‘screen-w’, if the terminal is wide (132 cols or more)). If even this entry cannot be found, ‘vt100’ is used as a substitute.

The idea is that if you have a terminal which doesn’t support an important feature (e.g. delete char or clear to EOS) you can build a new termcap/terminfo entry for screen (named ‘screen.dumbterm’) in which this capability has been disabled. If this entry is installed on your machines you are able to do a rlogin and still keep the correct termcap/terminfo entry. The terminal name is put in the $TERM variable of all new windows. screen also sets the $TERMCAP variable reflecting the capabilities of the virtual terminal emulated. Furthermore, the variable $WINDOW is set to the window number of each window.

The actual set of capabilities supported by the virtual terminal depends on the capabilities supported by the physical terminal. If, for instance, the physical terminal does not support underscore mode, screen does not put the ‘us’ and ‘ue’ capabilities into the window’s $TERMCAP variable, accordingly. However, a minimum number of capabilities must be supported by a terminal in order to run screen; namely scrolling, clear screen, and direct cursor addressing (in addition, screen does not run on hardcopy terminals or on terminals that over-strike).

Also, you can customize the $TERMCAP value used by screen by using the termcap command, or by defining the variable $SCREENCAP prior to startup. When the latter defined, its value will be copied verbatim into each window’s $TERMCAP variable. This can either be the full terminal definition, or a filename where the terminal ‘screen’ (and/or ‘screen-w’) is defined.

Note that screen honors the terminfo command if the system uses the terminfo database rather than termcap. On such machines the $TERMCAP variable has no effect and you must use the dumptermcap command (see Section 16.2 [Dump Termcap], page 80) and the tic program to generate terminfo entries for screen windows.

When the boolean ‘G0’ capability is present in the termcap entry for the terminal on which screen has been called, the terminal emulation of screen supports multiple character sets. This allows an application to make use of, for instance, the VT100 graphics character set or national character sets. The following control functions from ISO 2022 are supported:

‘lock shift G0’ (‘SI’), ‘lock shift G1’ (‘SO’), ‘lock shift G2’, ‘lock shift G3’, ‘single

shift G2’, and ‘single shift G3’. When a virtual terminal is created or reset, the ASCII character set is designated as ‘G0’ through ‘G3’. When the ‘G0’ capability is present, screen evaluates the capabilities ‘S0’, ‘E0’, and ‘C0’ if present. ‘S0’ is the sequence the terminal uses to enable and start the graphics character set rather than ‘SI’. ‘E0’ is the corresponding replacement for ‘SO’. ‘C0’ gives a character by character translation string that is used during semi-graphics mode. This string is built like the ‘acsc’ terminfo capability.

When the ‘po’ and ‘pf’ capabilities are present in the terminal’s termcap entry, applications running in a screen window can send output to the printer port of the terminal. This allows a user to have an application in one window sending output to a printer connected to the terminal, while all other windows are still active (the printer port is enabled and disabled again for each chunk of output). As a side-effect, programs running in different windows can send output to the printer simultaneously. Data sent to the printer is not displayed in the window. The info command displays a line starting with ‘PRIN’ while the printer is active.

Some capabilities are only put into the $TERMCAP variable of the virtual terminal if they can be efficiently implemented by the physical terminal. For instance, ‘dl’ (delete line) is only put into the $TERMCAP variable if the terminal supports either delete line itself or scrolling regions. Note that this may provoke confusion, when the session is reattached on a different terminal, as the value of $TERMCAP cannot be modified by parent processes. You can force screen to include all capabilities in $TERMCAP with the ‘-a’ command-line option (see Chapter 3 [Invoking Screen], page 5).

The "alternate screen" capability is not enabled by default. Set the altscreen .screenrc command to enable it.

16.2 Write out the window’s termcap entry

[Command] dumptermcap
     (C-a .)

Write the termcap entry for the virtual terminal optimized for the currently active window to the file .termcap in the user’s $HOME/.screen directory (or wherever screen stores its sockets. see Chapter 23 [Files], page 105). This termcap entry is identical to the value of the environment variable $TERMCAP that is set up by screen for each window. For terminfo based systems you will need to run a converter like captoinfo and then compile the entry with tic.

16.3 The termcap command

[Command] termcap term terminal-tweaks [window-tweaks]

[Command] terminfo term terminal-tweaks [window-tweaks]

[Command] termcapinfo term terminal-tweaks [window-tweaks]

(none)

Use this command to modify your terminal’s termcap entry without going through all the hassles involved in creating a custom termcap entry. Plus, you can optionally customize the termcap generated for the windows. You have to place these commands in one of the screenrc startup files, as they are meaningless once the terminal emulator is booted. If your system uses the terminfo database rather than termcap, screen will understand the terminfo command, which has the same effects as the termcap command. Two separate commands are provided, as there are subtle syntactic differences, e.g. when parameter interpolation (using ‘%’) is required. Note that the termcap names of the capabilities should also be used with the terminfo command. In many cases, where the arguments are valid in both terminfo and termcap syntax, you can use the command termcapinfo, which is just a shorthand for a pair of termcap and terminfo commands with identical arguments.

The first argument specifies which terminal(s) should be affected by this definition. You can specify multiple terminal names by separating them with ‘|’s. Use ‘*’ to match all terminals and ‘vt*’ to match all terminals that begin with ‘vt’.

Each tweak argument contains one or more termcap defines (separated by ‘:’s) to be inserted at the start of the appropriate termcap entry, enhancing it or overriding existing values. The first tweak modifies your terminal’s termcap, and contains definitions that your terminal uses to perform certain functions. Specify a null string to leave this unchanged (e.g. ""). The second (optional) tweak modifies all the window termcaps, and should contain definitions that screen understands (see Chapter 11 [Virtual Terminal], page 51).

16.4 Termcap Examples

Some examples:

     termcap xterm*  xn:hs@

Informs screen that all terminals that begin with ‘xterm’ have firm auto-margins that allow the last position on the screen to be updated (xn), but they don’t really have a status line (no ’hs’ – append ‘@’ to turn entries off). Note that we assume ‘xn’ for all terminal names that start with ‘vt’, but only if you don’t specify a termcap command for that terminal.

     termcap vt*  xn
     termcap vt102|vt220  Z0=\E[?3h:Z1=\E[?3l

Specifies the firm-margined ‘xn’ capability for all terminals that begin with ‘vt’, and the second line will also add the escape-sequences to switch into (Z0) and back out of (Z1) 132-character-per-line mode if this is a VT102 or VT220. (You must specify Z0 and Z1 in your termcap to use the width-changing commands.)

     termcap vt100  ""  l0=PF1:l1=PF2:l2=PF3:l3=PF4

This leaves your vt100 termcap alone and adds the function key labels to each window’s termcap entry.

     termcap h19|z19  am@:im=\E@:ei=\EO  dc=\E[P

Takes a h19 or z19 termcap and turns off auto-margins (am@) and enables the insert mode (im) and end-insert (ei) capabilities (the ‘@’ in the ‘im’ string is after the ‘=’, so it is part of the string). Having the ‘im’ and ‘ei’ definitions put into your terminal’s termcap will cause screen to automatically advertise the character-insert capability in each window’s termcap. Each window will also get the delete-character capability (dc) added to its termcap, which screen will translate into a line-update for the terminal (we’re pretending it doesn’t support character deletion).

If you would like to fully specify each window’s termcap entry, you should instead set the $SCREENCAP variable prior to running screen. See Chapter 11 [Virtual Terminal], page 51, for the details of the screen terminal emulation. See Chapter 16 [Termcap], page 79, for more information on termcap definitions.

16.5 Special Terminal Capabilities

The following table describes all terminal capabilities that are recognized by screen and are not in the termcap manual (see Chapter 16 [Termcap], page 79). You can place these capabilities in your termcap entries (in /etc/termcap) or use them with the commands termcap, terminfo and termcapinfo in your screenrc files. It is often not possible to place these capabilities in the terminfo database.

‘LP’ (bool)

Terminal has VT100 style margins (‘magic margins’). Note that this capability is obsolete — screen now uses the standard ‘xn’ instead.

‘Z0’ (str)

Change width to 132 columns.

‘Z1’ (str)

Change width to 80 columns.

‘WS’ (str)

Resize display. This capability has the desired width and height as arguments. SunView(tm) example: ‘\E[8;%d;%dt’.

‘NF’ (bool)

Terminal doesn’t need flow control. Send ^S and ^Q direct to the application. Same as flow off. The opposite of this capability is ‘nx’.

‘G0’ (bool)

Terminal can deal with ISO 2022 font selection sequences.

‘S0’ (str)

Switch charset ‘G0’ to the specified charset. Default is ‘\E(%.’.

‘E0’ (str)

Switch charset ‘G0’ back to standard charset. Default is ‘\E(B’.

‘C0’ (str)

Use the string as a conversion table for font 0. See the ‘ac’ capability for more details.

‘CS’ (str)

Switch cursor-keys to application mode.

‘CE’ (str)

Switch cursor-keys to cursor mode.

‘AN’ (bool)

Enable autonuke for displays of this terminal type. (see Section 16.6 [Autonuke], page 83).

‘OL’ (num)

Set the output buffer limit. See the ‘obuflimit’ command (see Section 16.7 [Obuflimit], page 83) for more details.

‘KJ’ (str)

Set the encoding of the terminal. See the ‘encoding’ command (see Section 11.11 [Character Processing], page 62) for valid encodings.

‘AF’ (str)

Change character foreground color in an ANSI conform way. This capability will almost always be set to ‘\E[3%dm’ (‘\E[3%p1%dm’ on terminfo machines).

‘AB’ (str)

Same as ‘AF’, but change background color.

‘AX’ (bool)

Does understand ANSI set default fg/bg color (‘\E[39m / \E[49m’).

‘XC’ (str)

Describe a translation of characters to strings depending on the current font. (see Section 16.8 [Character Translation], page 84).

‘XT’ (bool)

Terminal understands special xterm sequences (OSC, mouse tracking).

‘C8’ (bool)

Terminal needs bold to display high-intensity colors (e.g. Eterm).

‘TF’ (bool)

Add missing capabilities to the termcap/info entry. (Set by default).

16.6 Autonuke

[Command] autonuke state

(none) Sets whether a clear screen sequence should nuke all the output that has not been written to the terminal. See Section 16.7 [Obuflimit], page 83. This property is set per display, not per window.

[Command] defautonuke state

(none) Same as the autonuke command except that the default setting for new displays is also changed. Initial setting is off. Note that you can use the special AN terminal capability if you want to have a terminal type dependent setting.

16.7 Obuflimit

[Command] obuflimit [limit]

(none) If the output buffer contains more bytes than the specified limit, no more data will be read from the windows. The default value is 256. If you have a fast display (like xterm), you can set it to some higher value. If no argument is specified, the current setting is displayed. This property is set per display, not per window.

[Command] defobuflimit limit

(none)

Same as the obuflimit command except that the default setting for new displays is also changed. Initial setting is 256 bytes. Note that you can use the special OL terminal capability if you want to have a terminal type dependent limit.

16.8 Character Translation

Screen has a powerful mechanism to translate characters to arbitrary strings depending on the current font and terminal type. Use this feature if you want to work with a common standard character set (say ISO8851-latin1) even on terminals that scatter the more unusual characters over several national language font pages.

Syntax:

         XC=<charset-mapping>{,,<charset-mapping>}
         <charset-mapping> := <designator><template>{,<mapping>}
         <mapping> := <char-to-be-mapped><template-arg>

The things in braces may be repeated any number of times.

A <charset-mapping> tells screen how to map characters in font <designator> (‘B’: Ascii, ‘A’: UK, ‘K’: german, etc.) to strings. Every <mapping> describes to what string a single character will be translated. A template mechanism is used, as most of the time the codes have a lot in common (for example strings to switch to and from another charset). Each occurrence of ‘%’ in <template> gets substituted with the template-arg specified together with the character. If your strings are not similar at all, then use ‘%’ as a template and place the full string in <template-arg>. A quoting mechanism was added to make it possible to use a real ‘%’. The ‘\’ character quotes the special characters ‘\’, ‘%’, and ‘,’.

Here is an example:

         termcap hp700 'XC=B\E(K%\E(B,\304[,\326\\\\,\334]'

This tells screen, how to translate ISOlatin1 (charset ‘B’) upper case umlaut characters on a hp700 terminal that has a German charset. ‘\304’ gets translated to ‘\E(K[\E(B’ and so on. Note that this line gets parsed *three* times before the internal lookup table is built, therefore a lot of quoting is needed to create a single ‘\’.

Another extension was added to allow more emulation: If a mapping translates the unquoted ‘%’ char, it will be sent to the terminal whenever screen switches to the corresponding <designator>. In this special case the template is assumed to be just ‘%’ because the charset switch sequence and the character mappings normally haven’t much in common.

This example shows one use of the extension:

         termcap xterm 'XC=K%,%\E(B,[\304,\\\\\326,]\334'

Here, a part of the German (‘K’) charset is emulated on an xterm. If screen has to change to the ‘K’ charset, ‘\E(B’ will be sent to the terminal, i.e. the ASCII charset is used instead. The template is just ‘%’, so the mapping is straightforward: ‘[’ to ‘\304’, ‘\’ to ‘\326’, and ‘]’ to ‘\334’.

17 The Message Line

Screen displays informational messages and other diagnostics in a message line at the bottom of the screen. If your terminal has a status line defined in its termcap, screen will use this for displaying its messages, otherwise the last line of the screen will be temporarily overwritten and output will be momentarily interrupted. The message line is automatically removed after a few seconds delay, but it can also be removed early (on terminals without a status line) by beginning to type.

17.1 Using the message line from your program

The message line facility can be used by an application running in the current window by means of the ANSI Privacy message control sequence. For instance, from within the shell, try something like:

echo "<ESC>^Hello world from window $WINDOW<ESC>\"

where ‘<ESC>’ is ASCII ESC and the ‘^’ that follows it is a literal caret or up-arrow.

17.2 Hardware Status Line

[Command] hardstatus [state]

[Command] hardstatus [always]firstline|lastline|message|ignore

[string]

[Command] hardstatus string [string]

(none) This command configures the use and emulation of the terminal’s hardstatus line. The first form toggles whether screen will use the hardware status line to display messages. If the flag is set to ‘off’, these messages are overlaid in reverse video mode at the display line. The default setting is ‘on’. The second form tells screen what to do if the terminal doesn’t have a hardstatus line (i.e. the termcap/terminfo capabilities "hs", "ts", "fs" and "ds" are not set). If the type firstline/lastline is used, screen will reserve the first/last line of the display for the hardstatus. message uses screen’s message mechanism and ignore tells screen never to display the hardstatus. If you prepend the word always to the type (e.g., alwayslastline), screen will use the type even if the terminal supports a hardstatus line. The third form specifies the contents of the hardstatus line. %h is used as default string, i.e., the stored hardstatus of the current window (settable via ‘ESC]0;^G’ or ‘ESC_\\’) is displayed. You can customize this to any string you like including string escapes (see Chapter 21 [String Escapes], page 99). If you leave out the argument string, the current string is displayed. You can mix the second and third form by providing the string as additional argument.

17.3 Display Last Message

[Command] lastmsg

(C-a m, C-a C-m)

Repeat the last message displayed in the message line. Useful if you’re typing when a message appears, because (unless your terminal has a hardware status line) the message goes away when you press a key.

17.4 Message Wait

[Command] msgminwait sec

(none) Defines the time screen delays a new message when another is currently displayed. Defaults to 1 second.

[Command] msgwait sec

(none) Defines the time a message is displayed, if screen is not disturbed by other activity. Defaults to 5 seconds.

18 Logging

This section describes the commands for keeping a record of your session.

18.1 hardcopy

[Command] hardcopy [-h] [file]

(C-a h)

Writes out the currently displayed image to the file file, or, if no filename is specified, to hardcopy.n in the default directory, where n is the number of the current window. This either appends or overwrites the file if it exists, as determined by the hardcopy_append command. If the option -h is specified, dump also the contents of the scrollback buffer.

[Command] hardcopy_append state

(none) If set to ‘on’, screen will append to the hardcopy.n files created by the command hardcopy; otherwise, these files are overwritten each time.

[Command] hardcopydir directory

(none) Defines a directory where hardcopy files will be placed. If unset, hardcopys are dumped in screen’s current working directory.

18.2 log

[Command] deflog state

(none) Same as the log command except that the default setting for new windows is changed. Initial setting is ‘off’.

[Command] log [state]

(C-a H)

Begins/ends logging of the current window to the file screenlog.n in the window’s default directory, where n is the number of the current window. This filename can be changed with the ‘logfile’ command. If no parameter is given, the logging state is toggled. The session log is appended to the previous contents of the file if it already exists. The current contents and the contents of the scrollback history are not included in the session log. Default is ‘off’.

[Command] logfile filename

[Command] logfile flush secs

(none) Defines the name the log files will get. The default is ‘screenlog.%n’. The second form changes the number of seconds screen will wait before flushing the logfile buffer to the file-system. The default value is 10 seconds.

[Command] logtstamp [state]

[Command] logtstamp after secs
[Command] logtstamp string string

(none) This command controls logfile time-stamp mechanism of screen. If time-stamps are turned ‘on’, screen adds a string containing the current time to the logfile after two minutes of inactivity. When output continues and more than another two minutes have passed, a second time-stamp is added to document the restart of the output. You can change this timeout with the second form of the command. The third form is used for customizing the time-stamp string (‘-- %n:%t -- time-stamp -- %M/%d/%y %c:%s --\n’ by default).

19 Startup

This section describes commands which are only useful in the .screenrc file, for use at startup.

19.1 echo

[Command] echo [‘-n’] message

(none) The echo command may be used to annoy screen users with a ’message of the day’. Typically installed in a global screenrc. The option ‘-n’ may be used to suppress the line feed. See also sleep. Echo is also useful for online checking of environment variables.

19.2 sleep

[Command] sleep num

(none) This command will pause the execution of a .screenrc file for num seconds. Keyboard activity will end the sleep. It may be used to give users a chance to read the messages output by echo.

19.3 Startup Message

[Command] startup_message state

(none) Select whether you want to see the copyright notice during startup. Default is ‘on’, as you probably noticed.

20 Miscellaneous commands

The commands described here do not fit well under any of the other categories.

20.1 At

[Command] at [identifier][#|*|%] command [args]

(none) Execute a command at other displays or windows as if it had been entered there. At changes the context (the ‘current window’ or ‘current display’ setting) of the command. If the first parameter describes a non-unique context, the command will be executed multiple times. If the first parameter is of the form ‘identifier*’ then identifier is matched against user names. The command is executed once for each display of the selected user(s). If the first parameter is of the form ‘identifier%’ identifier is matched against displays. Displays are named after the ttys they attach. The prefix ‘/dev/’ or ‘/dev/tty’ may be omitted from the identifier. If identifier has a # or nothing appended it is matched against window numbers and titles. Omitting an identifier in front of the #, * or % character selects all users, displays or windows because a prefix-match is performed. Note that on the affected display(s) a short message will describe what happened. Note that the # character works as a comment introducer when it is preceded by whitespace. This can be escaped by prefixing # with a \. Permission is checked for the initiator of the at command, not for the owners of the affected display(s). Caveat: When matching against windows, the command is executed at least once per window. Commands that change the internal arrangement of windows (like other) may be called again. In shared windows the command will be repeated for each attached display. Beware, when issuing toggle commands like login! Some commands (e.g. \*Qprocess) require that a display is associated with the target windows. These commands may not work correctly under at looping over windows.

20.2 Break

[Command] break [duration]

(C-a b, C-a C-b)

Send a break signal for duration*0.25 seconds to this window. For non-Posix systems the time interval is rounded up to full seconds. Most useful if a character device is attached to the window rather than a shell process (see Section 6.6 [Window Types], page 29). The maximum duration of a break signal is limited to 15 seconds.

[Command] pow_break
(C-a B)

Reopen the window’s terminal line and send a break condition.

[Command] breaktype [tcsendbreak|TIOCSBRK|TCSBRK]

(none) Choose one of the available methods of generating a break signal for terminal devices. This command should affect the current window only. But it still behaves identical to defbreaktype. This will be changed in the future. Calling breaktype with no parameter displays the break setting for the current window.

[Command] defbreaktype [tcsendbreak|TIOCSBRK|TCSBRK]

(none) Choose one of the available methods of generating a break signal for terminal devices opened afterwards. The preferred methods are tcsendbreak and TIOCSBRK. The third, TCSBRK, blocks the complete screen session for the duration of the break, but it may be the only way to generate long breaks. tcsendbreak and TIOCSBRK may or may not produce long breaks with spikes (e.g. 4 per second). This is not only system dependent, this also differs between serial board drivers. Calling defbreaktype with no parameter displays the current setting.

20.3 Bumpleft

[Command] bumpleft

(none) Exchanges windows positions on window list, with window having lower number (left to current one).

20.4 Bumpright

[Command] bumpright

(none) Exchanges windows positions on window list, with window having bigger number (right to current one).

20.5 Collapse

[Command] collapse

(none) Changes windows numbers, so there is no gaps between them.

20.6 Debug

[Command] debug [on|off]

(none) Turns runtime debugging on or off. If screen has been compiled with option -DDEBUG debugging is available and is turned on per default. Note that this command only affects debugging output from the main ‘SCREEN’ process correctly. Debug output from attacher processes can only be turned off once and forever.

20.7 License

[Command] license

(C-a ,)

Display the disclaimer page. This is done whenever screen is started without options, which should be often enough.

20.8 Nethack

[Command] nethack state

(none) Changes the kind of error messages used by screen. When you are familiar with the game nethack, you may enjoy the nethack-style messages which will often blur the facts a little, but are much funnier to read. Anyway, standard messages often tend to be unclear as well. This option is only available if screen was compiled with the NETHACK flag defined (see Chapter 26 [Installation], page 111). The default setting is then determined by the presence of the environment variable $NETHACKOPTIONS and the file ~/.nethackrc - if either one is present, the default is on.

20.9 Nonblock

[Command] nonblock [state|numsecs]

Tell screen how to deal with user interfaces (displays) that cease to accept output. This can happen if a user presses ^S or a TCP/modem connection gets cut but no hangup is received. If nonblock is off (this is the default) screen waits until the display restarts to accept the output. If nonblock is on, screen waits until the timeout is reached (on is treated as 1s). If the display still doesn’t receive characters, screen will consider it “blocked” and stop sending characters to it. If at some time it restarts to accept characters, screen will unblock the display and redisplay the updated window contents.

[Command] defnonblock state|numsecs

Same as the nonblock command except that the default setting for displays is changed. Initial setting is off.

20.10 Number

[Command] number [[+|-]n]

(C-a N)

Change the current window’s number. If the given number n is already used by another window, both windows exchange their numbers. If no argument is specified, the current window number (and title) is shown. Using either a plus (‘+’) or minus (‘-’) will change the window’s number by the relative amount specified.

20.11 Time

[Command] time [string]

(C-a t, C-a C-t)

Uses the message line to display the time of day, the host name, and the load averages over 1, 5, and 15 minutes (if this is available on your system). For window-specific information use info (see Section 11.6 [Info], page 60). If a string is specified, it changes the format of the time report like it is described in the string escapes chapter (see Chapter 21 [String Escapes], page 99). Screen uses a default of ‘%c:%s %M %d

%H%? %l%?’.

20.12 Verbose

[Command] verbose [on|off]

If verbose is switched on, the command name is echoed, whenever a window is created (or resurrected from zombie state). Default is off. Without a parameter, the current setting is shown.

20.13 Version

[Command] version

(C-a v)

Display the version and modification date in the message line.

20.14 Zombie

[Command] zombie [keys [onerror] ]

[Command] zombie_timeout [seconds]

(none) Per default windows are removed from the window list as soon as the windows process (e.g. shell) exits. When a string of two keys is specified to the zombie command, ‘dead’ windows will remain in the list. The kill command may be used to remove the window. Pressing the first key in the dead window has the same effect. Pressing the second key, however, screen will attempt to resurrect the window. The process that was initially running in the window will be launched again. Calling zombie without parameters will clear the zombie setting, thus making windows disappear when the process terminates. As the zombie setting is affected globally for all windows, this command should probably be called defzombie, but it isn’t. Optionally you can put the word onerror after the keys. This will cause screen to monitor exit status of the process running in the window. If it exits normally (’0’), the window disappears. Any other exit value causes the window to become a zombie. Additionally the zombie_timeout command exists. If a window is declared “dead”, screen will automatically try to resurrect the window after the timeout. It only works if zombie keys are defined via zombie command.

20.15 Printcmd

[Command] printcmd [cmd]

(none) If cmd is not an empty string, screen will not use the terminal capabilities po/pf for printing if it detects an ansi print sequence ESC [ 5 i, but pipe the output into cmd. This should normally be a command like ‘lpr’ or ‘cat > /tmp/scrprint’. Printcmd without an argument displays the current setting. The ansi sequence ESC \ ends printing and closes the pipe. Warning: Be careful with this command! If other user have write access to your terminal, they will be able to fire off print commands.

20.16 Rendition

[Command] rendition bell | monitor | silence | so attr [color]

(none) Change the way screen renders the titles of windows that have monitor or bell flags set in caption or hardstatus or windowlist. See the chapter about string escapes (see Chapter 21 [String Escapes], page 99) for the syntax of the modifiers. The default for monitor is currently ‘=b’ (bold, active colors), for bell ‘=ub’ (underline, bold and active colors), and for silence ‘=u’.

20.17 Sorendition

[Command] sorendition [attr [color]]

(none) This command has been deprecated. Use rendition so instead.

20.18 Attrcolor

[Command] attrcolor attrib [attribute/color-modifier]

(none) This command can be used to highlight attributes by changing the color of the text. If the attribute attrib is in use, the specified attribute/color modifier is also applied. If no modifier is given, the current one is deleted. See the chapter about string escapes (see Chapter 21 [String Escapes], page 99) for the syntax of the modifier. Screen understands two pseudo-attributes, i stands for high-intensity foreground color and I for high-intensity background color. Examples:

attrcolor b "R"

Change the color to bright red if bold text is to be printed.

attrcolor u "-u b"

Use blue text instead of underline.

attrcolor b ".I"

Use bright colors for bold text. Most terminal emulators do this already.

attrcolor i "+b"

Make bright colored text also bold.

20.19 Setsid

[Command] setsid state

(none) Normally screen uses different sessions and process groups for the windows. If setsid is turned off, this is not done anymore and all windows will be in the same process group as the screen backend process. This also breaks job-control, so be careful. The default is on, of course. This command is probably useful only in rare circumstances.

20.20 Eval

[Command] eval command1 [command2 ...]

(none) Parses and executes each argument as separate command.

20.21 Maxwin

[Command] maxwin n

(none) Set the maximum window number screen will create. Doesn’t affect already existing windows. The number can be increased only when there are no existing windows.

20.22 Backtick

[Command] backtick id lifespan autorefresh command [args]
[Command] backtick id

(none) Program the backtick command with the numerical id id. The output of such a command is used for substitution of the %` string escape (see Chapter 21 [String Escapes], page 99). The specified lifespan is the number of seconds the output is considered valid. After this time, the command is run again if a corresponding string escape is encountered. The autorefresh parameter triggers an automatic refresh for caption and hardstatus strings after the specified number of seconds. Only the last line of output is used for substitution. If both the lifespan and the autorefresh parameters are zero, the backtick program is expected to stay in the background and generate output once in a while. In this case, the command is executed right away and screen stores the last line of output. If a new line gets printed screen will automatically refresh the hardstatus or the captions. The second form of the command deletes the backtick command with the numerical id id.

20.23 Screen Saver

[Command] idle [timeout [cmd args]]

(none) Sets a command that is run after the specified number of seconds inactivity is reached. This command will normally be the blanker command to create a screen blanker, but it can be any screen command. If no command is specified, only the timeout is set. A timeout of zero (ot the special timeout off) disables the timer. If no arguments are given, the current settings are displayed.

[Command] blanker

(none) Activate the screen blanker. First the screen is cleared. If no blanker program is defined, the cursor is turned off, otherwise, the program is started and it’s output is written to the screen. The screen blanker is killed with the first keypress, the read key is discarded. This command is normally used together with the idle command.

[Command] blankerprg [program args]

Defines a blanker program. Disables the blanker program if an empty argument is given. Shows the currently set blanker program if no arguments are given.

20.24 Zmodem

[Command] zmodem [off|auto|catch|pass]

[Command] zmodem sendcmd [string]

[Command] zmodem recvcmd [string]

(none) Define zmodem support for screen. Screen understands two different modes when it detects a zmodem request: pass and catch. If the mode is set to pass, screen will relay all data to the attacher until the end of the transmission is reached. In catch mode screen acts as a zmodem endpoint and starts the corresponding rz/sz commands. If the mode is set to auto, screen will use catch if the window is a tty (e.g. a serial line), otherwise it will use pass. You can define the templates screen uses in catch mode via the second and the third form. Note also that this is an experimental feature.

21 String Escapes

Screen provides an escape mechanism to insert information like the current time into messages or file names. The escape character is % with one exception: inside of a window’s hardstatus ^% (^E) is used instead.

Here is the full list of supported escapes: % the escape character itself a either am or pm A either AM or PM c current time HH:MM in 24h format C current time HH:MM in 12h format d day number D weekday name E sets %? to true if the escape character has been pressed. f flags of the window. See Section 10.7 [Windows], page 49, for meanings of the

various flags.

F sets %? to true if the window has the focus h hardstatus of the window H hostname of the system l current load of the system m month number M month name n window number P sets %? to true if the current region is in copy/paste mode s seconds S session name t window title u all other users on this window w all window numbers and names. With - qualifier: up to the current window;

with + qualifier: starting with the window after the current one.

W all window numbers and names except the current one y last two digits of the year number Y full year number ? the part to the next %? is displayed only if a % escape inside the part expands

to a non-empty string : else part of %? = pad the string to the display’s width (like TeX’s hfill). If a number is specified,

pad to the percentage of the window’s width. A 0 qualifier tells screen to treat the number as absolute position. You can specify to pad relative to the last absolute pad position by adding a + qualifier or to pad relative to the right margin by using -. The padding truncates the string if the specified position lies before the current position. Add the L qualifier to change this.

< same as %= but just do truncation, do not fill with spaces > mark the current text position for the next truncation. When screen needs to

do truncation, it tries to do it in a way that the marked position gets moved to the specified percentage of the output area. (The area starts from the last absolute pad position and ends with the position specified by the truncation operator.) The L qualifier tells screen to mark the truncated parts with ‘...’.

{ attribute/color modifier string terminated by the next } ` Substitute with the output of a ‘backtick’ command. The length qualifier is

misused to identify one of the commands. See Section 20.22 [Backtick], page 96. The c and C escape may be qualified with a 0 to make screen use zero instead of space

as fill character. The n and = escapes understand a length qualifier (e.g. %3n), D and M can be prefixed with L to generate long names, w and W also show the window flags if L is given.

An attribute/color modifier is is used to change the attributes or the color settings. Its

format is ‘[attribute modifier] [color description]’. The attribute modifier must be

prefixed by a change type indicator if it can be confused with a color description. The following change types are known: + add the specified set to the current attributes

- remove the set from the current attributes ! invert the set in the current attributes = change the current attributes to the specified set

The attribute set can either be specified as a hexadecimal number or a combination of the following letters: d dim u underline b bold r reverse s standout B blinking

Colors are coded either as a hexadecimal number or two letters specifying the desired background and foreground color (in that order). The following colors are known: k black r red g green y yellow b blue m magenta c cyan w white d default color . leave color unchanged

The capitalized versions of the letter specify bright colors. You can also use the pseudocolor ‘i’ to set just the brightness and leave the color unchanged.

A one digit/letter color description is treated as foreground or background color dependent on the current attributes: if reverse mode is set, the background color is changed instead of the foreground color. If you don’t like this, prefix the color with a ‘.’. If you want the same behavior for two-letter color descriptions, also prefix them with a ‘.’.

As a special case, ‘%{-}’ restores the attributes and colors that were set before the last change was made (i.e. pops one level of the color-change stack). Examples: ‘G’ set color to bright green ‘+b r’ use bold red ‘= yd’ clear all attributes, write in default color on yellow background.

‘%-Lw%{= BW}%50>%n%f* %t%{-}%+Lw%<’

The available windows centered at the current win dow and truncated to the available width. The current window is displayed white on blue. This can be

          used with ‘hardstatus alwayslastline’.
‘%?%F%{.R.}%?%3n %t%? [%h]%?’

The window number and title and the window’s hardstatus, if one is set. Also use a red background if this is the active focus. Useful for ‘caption string’.

22 Environment Variables

COLUMNS Number of columns on the terminal (overrides termcap entry). HOME Directory in which to look for .screenrc. LINES Number of lines on the terminal (overrides termcap entry). LOCKPRG Screen lock program.

NETHACKOPTIONS

Turns on nethack option.

PATH Used for locating programs to run.

SCREENCAP

For customizing a terminal’s TERMCAP value.

SCREENDIR

Alternate socket directory.

SCREENRC Alternate user screenrc file. SHELL Default shell program for opening windows (default /bin/sh). See also shell

.screenrc command.

STY Alternate socket name. If screen is invoked, and the environment variable STY is set, then it creates only a window in the running screen session rather than starting a new session.

SYSSCREENRC

Alternate system screenrc file.

TERM Terminal name. TERMCAP Terminal description. WINDOW Window number of a window (at creation time).

23 Files Referenced

.../screen-4.?.??/etc/screenrc
.../screen-4.?.??/etc/etcscreenrc

Examples in the screen distribution package for private and global initialization files.

$SYSSCREENRC
/local/etc/screenrc

screen initialization commands

$SCREENRC
$HOME/.iscreenrc
$HOME/.screenrc

Read in after /local/etc/screenrc

$SCREENDIR/S-login
/local/screens/S-login

Socket directories (default)

/usr/tmp/screens/S-login

Alternate socket directories.

socket directory/.termcap

Written by the dumptermcap command

/usr/tmp/screens/screen-exchange or
/tmp/screen-exchange

screen interprocess communication buffer

hardcopy.[0-9]

Screen images created by the hardcopy command

screenlog.[0-9]

Output log files created by the log command

/usr/lib/terminfo/?/* or
/etc/termcap

Terminal capability databases

/etc/utmp

Login records

$LOCKPRG Program for locking the terminal.

24 Credits

Authors =======

Originally created by Oliver Laumann. For a long time maintained and developed by Juergen Weigert, Michael Schroeder, Micah Cowan and Sadrul Habib Chowdhury. This latest version was produced by Amadeusz Slawinski <amade@asmblr.net> and Alexander Naumov <alexander naumov@opensuse.org>. Contributors ============

Thomas Renninger <treen@suse.com>,
Axel Beckert <abe@deuxchevaux.org>,
Ken Beal <kbeal@@amber.ssd.csd.harris.com>,
Rudolf Koenig <rfkoenig@@informatik.uni-erlangen.de>,
Toerless Eckert <eckert@@informatik.uni-erlangen.de>,
Wayne Davison <davison@@borland.com>,
Patrick Wolfe <pat@@kai.com, kailand!pat>,
Bart Schaefer <schaefer@@cse.ogi.edu>,
Nathan Glasser <nathan@@brokaw.lcs.mit.edu>,
Larry W. Virden <lvirden@@cas.org>,
Howard Chu <hyc@@hanauma.jpl.nasa.gov>,
Tim MacKenzie <tym@@dibbler.cs.monash.edu.au>,
Markku Jarvinen <mta@@@{cc,cs,ee@}.tut.fi>,
Marc Boucher <marc@@CAM.ORG>,
Doug Siebert <dsiebert@@isca.uiowa.edu>,
Ken Stillson <stillson@@tsfsrv.mitre.org>,
Ian Frechett <frechett@@spot.Colorado.EDU>,
Brian Koehmstedt <bpk@@gnu.ai.mit.edu>,
Don Smith <djs6015@@ultb.isc.rit.edu>,
Frank van der Linden <vdlinden@@fwi.uva.nl>,
Martin Schweikert <schweik@@cpp.ob.open.de>,
David Vrona <dave@@sashimi.lcu.com>,
E. Tye McQueen <tye%spillman.UUCP@@uunet.uu.net>,
Matthew Green <mrg@@eterna.com.au>,
Christopher Williams <cgw@@pobox.com>,
Matt Mosley <mattm@@access.digex.net>,
Gregory Neil Shapiro <gshapiro@@wpi.WPI.EDU>,
Jason Merrill <jason@@jarthur.Claremont.EDU>,
Johannes Zellner <johannes@@zellner.org>,
Pablo Averbuj <pablo@@averbuj.com>.

Version =======

This manual describes version 4.6.2 of the screen program. Its roots are a merge of a custom version 2.3PR7 by Wayne Davison and several enhancements to Oliver Laumann’s version 2.0. Note that all versions numbered 2.x are copyright by Oliver Laumann.

See also See Section 25.3 [Availability], page 110.

25 Bugs

Just like any other significant piece of software, screen has a few bugs and missing features. Please send in a bug report if you have found a bug not mentioned here.

25.1 Known Bugs

• ‘dm’ (delete mode) and ‘xs’ are not handled correctly (they are ignored). ‘xn’ is treated as a magic-margin indicator.

• screen has no clue about double-high or double-wide characters. But this is the only area where vttest is allowed to fail.

• It is not possible to change the environment variable $TERMCAP when reattaching under a different terminal type.

• The support of terminfo based systems is very limited. Adding extra capabilities to $TERMCAP may not have any effects.

• screen does not make use of hardware tabs.

• screen must be installed setuid root on most systems in order to be able to correctly change the owner of the tty device file for each window. Special permission may also be required to write the file /etc/utmp.

• Entries in /etc/utmp are not removed when screen is killed with SIGKILL. This will cause some programs (like "w" or "rwho") to advertise that a user is logged on who really isn’t.

• screen may give a strange warning when your tty has no utmp entry.

• When the modem line was hung up, screen may not automatically detach (or quit) unless the device driver sends a HANGUP signal. To detach such a screen session use the -D or -d command line option.

• If a password is set, the command line options -d and -D still detach a session without asking.

• Both breaktype and defbreaktype change the break generating method used by all terminal devices. The first should change a window specific setting, where the latter should change only the default for new windows.

• When attaching to a multiuser session, the user’s .screenrc file is not sourced. Each users personal settings have to be included in the .screenrc file from which the session is booted, or have to be changed manually.

• A weird imagination is most useful to gain full advantage of all the features.

25.2 Reporting Bugs

If you find a bug in Screen, please send electronic mail to ‘screen-devel@gnu.org’. Include the version number of Screen which you are using. Also include in your message the hardware and operating system, the compiler used to compile, a description of the bug behavior, and the conditions that triggered the bug. Please recompile screen with the ‘-DDEBUG’ options enabled, reproduce the bug, and have a look at the debug output written to the directory /tmp/debug. If necessary quote suspect passages from the debug output and show the contents of your config.h if it matters.

25.3 Availability

Screen is available under the GNU copyleft.

The latest official release of screen available via anonymous ftp from ‘prep.ai.mit.edu’, ‘nic.funet.fi’ or any other GNU distribution site. The home site of screen is

‘ftp.uni-erlangen.de (131.188.3.71)’, in the directory pub/utilities/screen. The

subdirectory ‘private’ contains the latest beta testing release. If you want to help, send a note to screen-devel@gnu.org.

26 Installation

Since screen uses pseudo-ttys, the select system call, and UNIX-domain sockets/named pipes, it will not run under a system that does not include these features of 4.2 and 4.3 BSD UNIX.

26.1 Socket Directory

The socket directory defaults either to $HOME/.screen or simply to /tmp/screens or preferably to /usr/local/screens chosen at compile-time. If screen is installed setuid root, then the administrator should compile screen with an adequate (not NFS mounted) SOCKDIR. If screen is not running setuid-root, the user can specify any mode 700 directory in the environment variable $SCREENDIR.

26.2 Compiling Screen

To compile and install screen:

The screen package comes with a GNU Autoconf configuration script. Before you compile the package run

                               sh ./configure

This will create a config.h and Makefile for your machine. If configure fails for some reason, then look at the examples and comments found in the Makefile.in

and config.h.in templates. Rename config.status to config.status.machine

when you want to keep configuration data for multiple architectures. Running sh ./config.status.machine recreates your configuration significantly faster than rerunning

configure.

Read through the "User Configuration" section of config.h, and verify that it suits your needs. A comment near the top of this section explains why it’s best to install screen setuid to root. Check for the place for the global screenrc-file and for the socket directory. Check the compiler used in Makefile, the prefix path where to install screen. Then run

                                    make

If make fails to produce one of the files term.h, comm.h or tty.c, then use filename.x.dist instead. For additional information about installation of screen refer to the file INSTALLATION, coming with this package.

Concept Index

.

.screenrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

A

availability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

B

binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 bug report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

C

capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 command character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 command line options . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 command summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 compiling screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 control sequences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 copy and paste. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 customization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

E

environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 escape character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

F

files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 flow control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

I

input translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

K

key binding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

L

layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

M

marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 message line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 multiuser session . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

O

options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

R

regions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

S

screenrc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 scrollback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 socket directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 string escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

T

terminal capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

W

window groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 window types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Command Index

This is a list of all the commands supported by screen.

A

acladd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 aclchg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 acldel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 aclgrp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 aclumask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 activity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 addacl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 allpartial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 altscreen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 attrcolor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 autodetach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 autonuke . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

B

backtick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 bce . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 bell_msg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 bindkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 blanker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 blankerprg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 breaktype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 bufferfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 bumpleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 bumpright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

C

c1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 caption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 chacl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 chdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 cjkwidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 clear. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 collapse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 colon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 compacthist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 copy_reg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 crlf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

D

debug. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 defautonuke . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 defbce. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 defbreaktype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 defc1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 defcharset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 defencoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 defescape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 defflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 defgr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 defhstatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 deflog. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 deflogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 defmode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 defmonitor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 defmousetrack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 defnonblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 defobuflimit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 defscrollback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 defshell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 defsilence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 defslowpaste. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 defutf8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 defwrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 defwritelock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 detach. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 digraph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 dinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 displays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 dumptermcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

E

echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 escape. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

F

fit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 flow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 focus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 focusminsize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

G

gr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

116 Screen User’s Manual

H

hardcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 hardcopy_append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 hardcopydir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 hardstatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 hstatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

I

idle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ignorecase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 ins_reg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

K

kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

L

lastmsg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42, 43 license . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 lockscreen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 logfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 login. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 logtstamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

M

mapdefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 mapnotnext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 maptimeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 markkeys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 maxwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 meta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 monitor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 mousetrack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 msgminwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 msgwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 multiuser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

N

nethack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 nonblock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

O

obuflimit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 other. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

P

partial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 paste. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 pastefont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 pow_break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 pow_detach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 pow_detach_msg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 prev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 printcmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

Q

quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

R

readbuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 readreg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 redisplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 register . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 remove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 removebuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 rendition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 reset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 resize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

S

screen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 scrollback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 select. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 sessionname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 setenv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 setsid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 shelltitle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 silence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 silencewait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 sleep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 slowpaste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 sorendition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 split. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 startup_message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 stuff. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 su . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

Command Index 117

T

term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 termcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 termcapinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 terminfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

U

umask. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 unbindall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 utf8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

V

vbell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 vbell_msg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 vbellwait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

W

wall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 windowlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 writebuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 writelock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

X

xoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 xon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

Z

zmodem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 zombie. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 zombie_timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Keystroke Index

This is a list of the default key bindings.

The leading escape character (see Section 14.3 [Command Character], page 74) has been omitted from the key sequences, since it is the same for all bindings.

"_{" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31} ’_{' . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31} *_{* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35} ,_{, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92} ._{. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80} :_{: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9} <_{< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70} =_{= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70} >_{> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70} ?_{? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74} [_{[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65} ]_{] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68} _. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

\

\ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

{

{ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

}

} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

|

| . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

0

0...9. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

A

a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

B

b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 Backspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

120 Screen User’s Manual

C

c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 C-[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 C-] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 C-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 C-d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 C-g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 C-h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 C-i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 C-k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 C-l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 C-m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 C-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 C-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 C-x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C-z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

E

ESC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

F

f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

I

i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

K

k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 SPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 TAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

V

v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

W

w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

X

x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

Z

z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62